### v1

# SMART: Self-Generating and Self-Validating Multi-Dimensional Assessment for LLMs’ Mathematical Problem Solving

Abstract
Large Language Models have achieved remarkable results on a variety of mathematical benchmarks. However, concerns remain as to whether these successes reflect genuine mathematical reasoning or superficial pattern recognition. Common evaluation metrics, such as final answer accuracy, fail to disentangle the underlying competencies involved, offering limited diagnostic value. To address these limitations, we introduce SMART: a Self-Generating and Self-Validating Multi-Dimensional Assessment Framework. SMART decomposes mathematical problem solving into four distinct dimensions: understanding, reasoning, arithmetic, and reflection & refinement. Each dimension is evaluated independently through tailored tasks, enabling interpretable and fine-grained analysis of LLM behavior. Crucially, SMART integrates an automated self-generating and self-validating mechanism to produce and verify benchmark data, ensuring both scalability and reliability. We apply SMART to 21 state-of-the-art open- and closed-source LLMs, uncovering significant discrepancies in their abilities across different dimensions. Our findings demonstrate the inadequacy of final answer accuracy as a sole metric and motivate a new holistic metric to better capture true problem-solving capabilities. Code and benchmarks will be released upon acceptance.
SMART: Self-Generating and Self-Validating Multi-Dimensional Assessment for LLMs’ Mathematical Problem Solving
Yujie Hou, Ting Zhang, Mei Wang, Xuetao Ma, Hua Huangthanks:  Corresponding author: Hua Huang School of Artificial Intelligence, Beijing Normal University, Beijing, China {houyujie, maxuetao}@mail.bnu.edu.cn, {tingzhang,wangmei1, huahuang}@bnu.edu.cn
1 Introduction
Large language models (LLMs)  Achiam et al. (2023); Wei et al. (2022) have demonstrated impressive performance across various natural language processing tasks Singhal et al. (2023); Wang et al. (2024); Jiang et al. (2024); Ma et al. (2025) and are being increasingly integrated into real-world applications, such as education, scientific computing, and decision support OpenAI (2024); Guo et al. (2025). As their adoption accelerates, it is critical to rigorously evaluate their true capabilities.
Figure 1: Following Polya’s four-step framework for human mathematical problem solving—understanding the problem, devising a plan, carrying out the plan, and looking back—we define four corresponding evaluation dimensions and design specific tasks to systematically assess the capabilities of large language models.
While numerous benchmarks have been proposed to evaluate the mathematical reasoning capabilities of LLMs  Cobbe et al. (2021); Ling et al. (2017); Patel et al. (2021); Miao et al. (2020); Koncel-Kedziorski et al. (2016), many suffer from a significant limitation: they provide insufficient interpretability regarding actual mathematical proficiency. The predominant focus on final answer accuracy is particularly problematic, as this metric reflects an aggregate outcome influenced by multiple, intertwined cognitive processes. As a result, existing benchmarks fail to effectively evaluate the distinct components of the mathematical problem-solving process Cobbe et al. (2021); Ling et al. (2017); Patel et al. (2021); Miao et al. (2020); Koncel-Kedziorski et al. (2016).
Another critical deficiency stems from the static nature of most existing benchmarks. Such static evaluation sets are highly susceptible to data contamination, potentially leading to inflated performance scores and skewed assessments of LLM capabilities Oren et al. (2023); Zhu et al. (2023, 2024); Li et al. (2024a). Furthermore, as LLMs rapidly advance, they can quickly reach performance saturation on these fixed datasets. Consequently, static benchmarks may no longer accurately reflect the true capabilities of more advanced models Fodor (2025); AI@Meta (2024).
Furthermore, constructing high-quality benchmark datasets is resource-intensive, requiring significant human labor and time for annotation Mirzadeh et al. (2024); Kurtic et al. (2024). When leveraging advanced LLMs like GPT-4o for data generationLi et al. (2024b); Zhu et al. (2024), ensuring the accuracy and consistency of the generated content remains a major challenge, as LLMs may produce errors and inconsistencies. Although human annotators are often employed to review and refine LLM-generated questions and answers Zhu et al. (2024); Zheng et al. (2023), this manual intervention is costly and inherently limits scalability.
Figure 2: An overview of the SMART framework for evaluating the mathematical problem-solving process. The SMART contains four distinct dimension. Each dimension is evaluated using dimension-specific task and metrics, ensuring a comprehensive assessment of the model’s problem-solving capabilities.
In light of these issues, we propose a novel evaluation framework: the Self-Generating and Self-Validating Multi-Dimensional Assessment Framework (SMART). Inspired by Polya’s problem-solving theory Polya (2014), SMART deconstructs the mathematical problem-solving process into four key evaluation dimensions: understanding, reasoning, arithmetic, and reflection & refinement. Each dimension is designed to evaluate a distinct facet of problem-solving progression (Fig. 1).
To implement this multi-dimensional assessment, SMART generates dimension-specific tasks with clearly defined outputs from seed problems. This structured decomposition allows SMART to enable the model to probe distinct cognitive components of mathematical problem solving. By isolating each dimension, SMART facilitates comprehensive evaluation and yields insights into the strengths and weaknesses of LLMs, informing targeted algorithmic enhancements. The resulting evaluation dataset, SMART-Bench, is built through an integrated self-generating and self-validating mechanism that not only programmatically produces dimension-specific data but also iteratively verifies its quality. This approach ensures both scalability and reliability in benchmark construction.
We evaluate 21 recently released open- and closed-source LLMs on SMART-Bench. Our experimental results reveal that models exhibit diverse performance patterns, demonstrating varied strengths and weaknesses across different dimensions. Furthermore, our investigation pinpoints key factors influencing performance across these distinct dimensions. Finally, through an in-depth analysis of the relationship between final answer accuracy and the scores of our decomposed dimensions, we propose a new metric designed to more accurately measure the true mathematical capabilities of LLMs. Our main contributions are as follows:
  1. $\bullet$
    We propose SMART, a self-generating and self-validating multi-dimensional assessment framework designed to break down the mathematical problem-solving process of LLMs into four key dimensions. This approach enables deeper insight into underlying cognitive processes beyond final answer accuracy.
  2. $\bullet$
    We introduce an integrated self-generating and self-validating mechanism that produces SMART-bench dataset. This significantly reduces reliance on human annotation while ensuring the benchmark’s scalability and reliability.
  3. $\bullet$
    We perform extensive evaluations revealing substantial disparities in the mathematical abilities of various LLMs, providing detailed, dimension-specific, and interpretable analyses. These findings offer a valuable foundation for future LLM development and optimization.
2 Related Work
Mathematic benchmark. Numerous mathematical benchmarks with varying levels of difficulty have been developed to explore the upper bound of LLMs’ mathematical capabilities. These benchmarks range from grade-school-level datasets Cobbe et al. (2021), to high-school-level datasets Hendrycks et al. (2021), and extend to expert-level datasets Glazer et al. (2024). Their scope covers a broad range of mathematical domains, including geometry, number theory, and real analysis. However, despite the increasing difficulty of these benchmarks, it remains unclear whether LLMs truly comprehend mathematical concepts or merely rely on pattern-matching to generate correct answers Mirzadeh et al. (2024). Furthermore, the widespread use of these benchmarks increase the risk of data contamination, potentially inflating performance evaluations Li et al. (2024a).
Dynamic evaluation. To address these concerns, recent studies have explored dynamic evaluation approaches Zhu et al. (2023, 2024), which generate adaptive test data based on pre-defined rewriting rules. For example, GSM-Plus Li et al. (2024b) introduces eight perturbation strategies, while GSM-Symbolic Mirzadeh et al. (2024) generates diverse problem variants using symbolic templates. These approaches aim to mitigate data leakage and improve robustness in assessments. Despite these advancements, current evaluation methodologies remain limited. They predominantly focus on final-answer accuracy, which fails to capture an LLM’s reasoning process in a systematic manner. Additionally, they do not explicitly assess the diverse cognitive skills required for mathematical problem-solving, nor do they provide fine-grained insights into the underlying logical reasoning steps.
These limitations underscore the need for a comprehensive, interpretable, and scalable evaluation framework that can decompose the problem-solving process, assess reasoning at multiple levels, and minimize reliance on human verification. To bridge this gap, we introduce SMART, a self-validating multi-dimensional assessment framework designed to systematically evaluate LLMs’ mathematical reasoning abilities while enhancing benchmark reliability and scalability.
3 Methodology
3.1 SMART Framework
Understanding. The initial step in mathematical problem-solving is to comprehend the problem. Accordingly, the first dimension in the SMART framework assesses the ability of LLMs to clearly and comprehensively understand a given problem before attempting a solution. To rigorously evaluate this dimension, we introduce a context extraction task. This task requires the LLM to meticulously identify and categorize various informational components embedded within the problem. As illustrated in Fig. 2 (Dimension 1, Task 1), the LLM is prompted to extract and structure key information into a predefined template. This template includes categories such as: problem scenario, goal, known and unknown quantities, relationships and constraints, and irrelevant information.
This structured context extraction task serves a dual purpose. At a surface level, it evaluates the LLM’s ability to summarize and extract salient information. More profoundly, it probes the depth of the LLM’s comprehension by requiring it to discern the roles and interconnections of different problem elements, thereby offering a nuanced measure of its problem understanding capabilities. This moves beyond simple keyword spotting to assess a more holistic grasp of the problem’s structure.
Reasoning. Subsequent to the comprehension phase, Polya’s problem-solving heuristic emphasizes devising a coherent plan. Aligning with this, the second dimension evaluates the reasoning ability of LLMs. This is achieved through a symbolic formalization task, wherein LLMs are required to solve mathematical problems via formal symbolic expressions. This approach compels the LLM to concentrate on the inherent logical structure of the problem and the intricate relationships between its constituent elements, rather than relying on direct numerical computation.
Beyond merely translating problem statements into symbolic form, this task demands the identification of solution-relevant equations and the construction of a logically consistent path to the answer. As illustrated in Fig. 2 (Dimension 2, Task 2), a highlighted symbolic assertion, absent from the original question, captures a key inferred relation, indicating the model’s ability to derive implicit knowledge. This dimension thus probes deeper reasoning by emphasizing inference and the construction of a complete logical model.
We adopt a standardized and widely used format for representing satisfiability modulo theories (SMT) problems Barrett et al. (2010). The SMT-LIB expressions produced by the models are subsequently verified using established SMT solvers such as Z3 De Moura and Bjørner (2008), which can automatically determine the satisfiability of logical formulas or compute concrete solutions. A reasoning plan expressed in SMT-LIB is considered correct if the solution derived by the solver consistently align with the ground truth of the original mathematical problem.
Arithmetic. The third dimension, Arithmetic, is dedicated to assessing an LLM’s proficiency in performing accurate numerical computations. This evaluation is conducted through a numeric calculation task that presents problems exclusively in terms of numerical values and variables, requiring the execution of fundamental arithmetic operations. For instance, as illustrated in Fig.  2 (Dimension 3, Task 3), the arithmetic questions are designed to contain only basic calculation operations and clearly defined variables.
This task design isolates arithmetic skills by minimizing confounding cognitive demands such as language comprehension or complex reasoning, thereby enabling a focused and accurate evaluation of the model’s computational proficiency.
Reflection & Refinement. The final stage in Polya’s problem-solving model, "looking back," involves reviewing the solution and improving its correctness. Recent literature has demonstrated that LLMs possess a promising capacity for self-feedback, and iterative refinement Madaan et al. (2023); Shinn et al. (2023); Guo et al. (2025); Huang et al. (2022). These studies suggest that prompting LLMs to scrutinize their own outputs or follow structured verification steps lead to improved accuracy and more robust reasoning. So, it is crucial to evaluate the self-critique capability of LLMs.
Accordingly, the fourth dimension of our SMART framework, Reflection & Refinement, aims to evaluate the LLM’s ability to critically review its problem-solving process and refine its answers. This is operationalized through an error correction task, an aspect largely overlooked in prior benchmarks. As depicted in Fig. 2 (Dimension 4, Task 4), LLMs are presented with a Chain-of-Thought (CoT) solution that contains deliberately introduced errors. We introduce six distinct categories of errors into the CoT, encompassing arithmetic number error, skipped step error, hallucinatory injection error, logical order error, redundant output error, and incorrect operator error (detailed in Fig. 18). The LLM is then tasked with first identifying any flaws within the provided CoT, then correcting these identified errors, and finally generating a refined answer based on the corrected reasoning. A critical design choice is that if an LLM fails to detect any errors in the provided CoT, it does not proceed to the subsequent refinement tasks for that particular problem instance.
3.2 The Self-Generating and Self-Validating Mechanism in SMART
A core innovation of SMART lies in its integrated self-generating and self-validating mechanism, which ensures both scalable benchmark construction and the reliability of SMART-Bench data.
3.2.1 Dataset Generation
To construct SMART-Bench, we begin by curating a diverse set of 2,000 seed questions, selected from seven widely used mathematical problem datasets: GSM8K Cobbe et al. (2021), SVAMP Patel et al. (2021), ASDiv Miao et al. (2020), AQuA Ling et al. (2017), MAWPS Koncel-Kedziorski et al. (2016), MATH Hendrycks et al. (2021), and the AIME 2024 competition Huggingface (2024). For each seed question, we automatically generate four dimension-specific variants, each aligned with a core dimension. The resulting dataset comprises 10,000 instances, 2,000 original questions and 8,000 generated variants. Details about generation procedures are provided in Appendix B.
3.2.2 Dataset Verification
For the understanding dimension, given the strong capability of LLMs in information extraction, we rely on a robust LLM to ensure correctness. In contrast, for reasoning and arithmetic questions, manual verification by human annotators is resource-intensive and limits scalability. To overcome this, SMART incorporates a self-validating mechanism based on a neuro-symbolic framework that automatically and precisely verifies generated question variants and their answers for dimensions suitable for symbolic representation. For example, to verify the correctness of a generated SMT-LIB expression for a seed question, we employ an SMT solver to compute the symbolic formula’s answer automatically. We then compare the solver’s output with the known seed answer, iterating this process until the generated SMT-LIB expression aligns with the expected result. Further details are provided in Appendix C. For the Reflection & Refinement dimension, ground truth is guaranteed since errors are intentionally produced.
To further assess the robustness of LLM capabilities, SMART employs a suite of rewriting rules to create augmented versions of the dimension-specific tasks. The specifics of these rewriting strategies and the results of associated experiments are detailed in Sec. 4.3 and 4.4. The self-validation methodology remains effective for ascertaining the correctness of these newly formulated, augmented questions, even when their answers differ from those of the original problems. This capability is crucial for maintaining the integrity of dynamically generated test instances. A comprehensive description of the procedures for the self-validation of augmented data is available in Appendix C.
4 Experiments
4.1 Experimental Setup
We evaluate 21 recent open-source and closed-source LLMs using our SMART evaluation framework. The open-source models include Phi4 Abdin et al. (2024), Gemma3 Team et al. (2025), GLM4 GLM et al. (2024), Qwen2.5 Team (2024), Qwen3 Yang et al. (2025), Llama3 AI@Meta (2024), Llama4 Meta (2025), and Mistral MistralAITeam (2024). The closed-source models assessed are GPT-4o Achiam et al. (2023), GPT-4.1 OpenAI (2025a), o4-mini, o3 OpenAI (2025b), DeepSeek-V3 Liu et al. (2024), DeepSeek-R1 Guo et al. (2025), Claude3.5 anthropic (2024), Claude3.7 anthropic (2025), and Gemini2.5 Google (2024). The temperature is set to 0 to ensure deterministic outputs. We employ an LLM-as-a-judge approach Zheng et al. (2023) to evaluate the understanding dimension. Tasks for the other dimensions primarily utilize accuracy-based metrics. For all dimension-specific tasks, we adopt a three-shot prompting strategy. Detailed descriptions of the evaluation metrics and prompt designs for each of the four SMART dimensions are provided in Appendix D.2.
Model Final Answer Understanding Reasoning Arithmetic Reflection&Refinement Reflection Refinement
ACC@Fi LLM@Un ACC@Re ACC@Ar ACC@RR ACC@R-t ACC@R-m
Open-source models
Phi4-14B 89.55 93.17 64.12 96.45 22.06 26.75 82.34
Gemma3-27B 54.62 94.47 40.81 39.94 24.23 32.55 74.35
GLM4-32B 54.13 94.39 55.92 36.62 22.85 28.50 88.57
Qwen2.5-72B 41.05 95.21 62.53 35.55 33.85 35.75 94.69
Qwen3-32B 93.75 93.19 86.14 98.50 54.75 56.70 96.52
Qwen3-30B-A3B 92.85 93.29 85.20 98.15 43.89 58.25 75.36
Llama3.1-8B 45.15 88.92 9.05 62.15 3.30 8.80 37.46
Llama3.3-70B 88.95 95.04 49.26 94.72 40.30 41.75 96.53
Llama4-Scout-17B-16E 91.24 94.48 62.65 97.11 29.04 49.5 58.67
Mistral-Small-3.1-24B 77.55 94.48 28.96 49.55 24.55 36.65 66.98
Mistral-Large-123B 62.25 94.75 48.95 43.74 43.12 50.61 81.62
Closed-source models
GPT-4o 57.65 95.32 64.45 45.64 42.60 45.30 94.04
GPT-4.1 59.31 95.41 73.29 45.71 59.90 63.80 93.89
o4-mini 93.75 95.22 90.72 98.25 55.05 64.61 85.22
o3 91.44 95.17 89.12 98.45 71.51 78.62 90.97
DeepSeek-V3 93.43 94.92 65.25 98.35 68.65 71.10 96.55
DeepSeek-R1 97.92 94.88 93.61 98.15 62.50 69.80 89.54
Claude3.5-Sonnet 92.61 94.69 80.85 97.61 33.05 65.85 50.19
Claude3.7-Sonnet 93.53 95.09 74.18 98.11 57.43 66.35 86.51
Gemini2.5-Flash-Preview 93.33 94.25 89.72 98.12 35.85 50.35 71.20
Gemini2.5-Pro-Preview 94.15 94.23 91.45 97.95 54.80 76.90 71.26
Table 1: The performance of open and closed-source models on SMART-Bench.
4.2 Performance on the SMART-Bench
The performance of the 21 evaluated LLMs on SMART-Bench is detailed in Tab. 1, which presents scores for final answer accuracy alongside metrics for each of the four dimensions.
Dimension specific-task results. Our results indicate that the evaluated models generally demonstrate a strong capacity for problem understanding, with most LLM@Un scores exceeding 90%. This suggests a general proficiency in grasping relevant information and interpreting problem statements. However, significant performance disparities emerge in the reasoning dimension, where ACC@Re scores range widely from 9.05% to 91.45%. A similar divergence is observed in the reflection task with the highest score (78.62%) being nearly nine times greater than the lowest (8.80%). These findings suggest that symbolic reasoning and error reflection capabilities represent critical bottlenecks, particularly for lower-performing models. In contrast, the arithmetic and refinement tasks appear relatively more tractable for most models, with top-performing LLMs achieving near-perfect scores. For instance, o3 attains an ACC@Ar of 98.45%, and DeepSeek-V3 achieves an ACC@R-m of 96.55%, underscoring their strengths in computational and corrective abilities.
Granular insights. Furthermore, the SMART-Bench framework reveals nuanced performance differences among LLMs that final answer accuracy can obscure. For example, while o4-mini and Claude3.7-Sonnet exhibit similar final answer accuracies (93.75% and 93.53%, respectively), o4-mini demonstrates markedly higher proficiency in the reasoning dimension (90.72%) compared to Claude3.7-Sonnet (74.18%). A similar pattern is observed when comparing o4-mini with DeepSeek-V3, demonstrating SMART-Bench’s capacity to reveal fine-grained performance gaps beyond final accuracy metrics. Additionally, consider o4-mini and GPT-4.1: they achieve comparable scores in the understanding dimension and the reflection & refinement dimension. However, o4-mini’s final answer accuracy (93.75%) substantially surpasses that of GPT-4.1 (59.31%). Our framework attributes this disparity primarily to GPT-4.1’s lower reasoning (73.29% vs. o4-mini’s 90.72%) and arithmetic (45.71% vs. o4-mini’s 98.25%) scores. Thus, the SMART framework facilitates a deeper analysis and interpretation of the underlying causes for performance differences.
Data contamination. Theoretically, LLMs should achieve higher accuracy on the arithmetic task compared to the final answer accuracy on the more complex seed questions, as the arithmetic tasks present simplified numerical computations in mathematical notation. However, our results reveal instances where models such as Gemma3-27B, GLM4-32B, GPT-4o, and GPT-4.1 exhibit lower ACC@Ar scores than their ACC@Fi scores. This counter-intuitive finding may suggest the presence of data contamination in the training sets of these models, where they might have memorized answers to full problems similar to those in standard benchmarks but struggle with the decontextualized arithmetic components presented in our SMART-Bench.
Figure 3: The performance across the varying difficulty settings for each SMART dimension.
4.3 How Does Task Difficulty Impact Different Dimensions of SMART?
To examine factors affecting model performance across each SMART dimension, we construct new sets of dimension-specific questions with varying difficulty levels and evaluate five leading closed-source LLMs on this dynamic test set. Details of the difficulty design are provided in Appendix D.3.
Fig. 3 presents performance trends across difficulty levels for each SMART dimension. As expected, increased task complexity generally leads to substantial performance degradation, validating the effectiveness of our difficulty calibration. Notably, GPT-4.1 and Claude3.7-Sonnet exhibit heightened sensitivity in the reasoning dimension, with ACC@Re scores dropping sharply from  90% to below 60% as reasoning steps increase. In contrast, the remaining models sustain ACC@Re scores above 80% even beyond six reasoning steps. In the arithmetic dimension, o4-mini maintains robust performance (ACC@Ar > 80%) even with nine-digit numbers (refers to the number of digits in scientific notation instead of the count of operands), whereas GPT-4.1’s accuracy drops below 10%. For reflection tasks, introducing only two error types into the CoT results in a sharp decline in detection accuracy across all models. None can reliably detect all errors when four or more distinct mistake types are present.
Model Perturbation Final Answer Understanding Reasoning Arithmetic
ACC@An PD LLM@Un PD ACC@Reason PD ACC@Ar PD
GPT-4.1 Seed question 59.31 / 95.41 / 73.29 / 45.71 /
+ Noise insertion 41.23 18.08$\downarrow$ 88.52 6.89$\downarrow$ 46.54 26.75$\downarrow$ 39.40 6.31$\downarrow$
+ Adding operation 28.45 30.86$\downarrow$ 91.32 4.09$\downarrow$ 57.53 15.76$\downarrow$ 20.64 25.07$\downarrow$
+ Numerical variation 24.81 34.50$\downarrow$ 92.83 2.58$\downarrow$ 52.81 20.48$\downarrow$ 18.34 27.37$\downarrow$
Claude-3.7 Seed question 93.53 / 95.09 / 74.18 / 98.11 /
+ Noise insertion 87.60 5.93$\downarrow$ 89.87 5.22$\downarrow$ 65.81 8.37$\downarrow$ 91.82 6.29$\downarrow$
+ Adding operation 76.21 17.32$\downarrow$ 92.34 2.75$\downarrow$ 54.38 19.80$\downarrow$ 67.43 24.39$\downarrow$
+ Numerical variation 62.74 30.79$\downarrow$ 93.17 1.92$\downarrow$ 58.92 15.26$\downarrow$ 58.29 39.82$\downarrow$
Table 2: The performance degradation of evaluation dimensions when three types of perturbations are added to the seed questions. PD refers to the performance drop.
4.4 How Do Fine-grained Dimensions Influence the Performance of Final Answer Accuracy?
Prior work has demonstrated that LLMs exhibit significant performance degradation in final answer accuracy when evaluated on question variations generated through perturbations Li et al. (2024b); Zhu et al. (2023); Li et al. (2024a). However, the specific contributions of different cognitive dimensions to this degradation remain underexplored. Therefore, we adapt three perturbation strategies from Li et al. (2024b) applying them to both the seed questions and their corresponding dimension-specific variants. This allows us to investigate which underlying capabilities are most influential in the observed drop in final answer accuracy under perturbation.
Tab. 2 shows that all evaluated dimensions exhibit noticeable performance drops (PD) under perturbations. When noise is introduced, the reasoning dimension is most affected for both GPT-4.1 (26.75%) and Claude 3.7-Sonnet (8.37%). In contrast, perturbations involving added operations or numerical modifications result in the largest performance drops in the arithmetic dimension for both models. These findings suggest that irrelevant information primarily disrupts reasoning capabilities, while alterations to operations or numerical values predominantly impair arithmetic proficiency. Such fine-grained dimension performance drop ultimately contribute to reduced final answer accuracy.
Figure 4: The confusion matrix of final answer and other dimensions. P means Positive, and N means Negative.
4.5 Is the Final Answer Accuracy Reliable for Measuring Mathematical Capability?
Given the impressive performance of LLMs and the potential for data contamination, there is a concern that models might solve problems correctly without possessing genuine underlying mathematical capability. To investigate this, we conduct experiments to compute confusion matrices comparing final answer correctness with performance on our SMART dimensions, as illustrated in Fig. 4. We posit that true mathematical capability is more accurately reflected by instances where LLMs correctly solve not only the original seed question but also simultaneously succeed in the corresponding reasoning and arithmetic dimension tasks.
Across all confusion matrices, the False Negative (FN) values are consistently non-zero. This indicates that LLMs can sometimes arrive at correct final answers through heuristic shortcuts or other opaque mechanisms, even when their intermediate reasoning or calculation processes are flawed. For example, GPT-4.1 exhibits a notable FN rate of 11.90% in final answer and arithmetic confusion matrix. Similarly, Claude3.7-Sonnet shows an FN rate of 22.95% in final answer and reasoning confusion matrix. These FN cases represent instances where final answer accuracy overestimates the model’s grasp of the intermediate steps.
Conversely, False Positive (FP) scores denote cases where a model successfully completes an intermediate task but ultimately yields an incorrect final answer. Except for GPT-4.1, most evaluated LLMs exhibit relatively low FP rates across various confusion matrices, indicating that accurate intermediate reasoning generally correlates with correct final outputs.
True Positive (TP) scores capture instances where a model not only produces the correct final answer but also performs intermediate reasoning and arithmetic correctly. We regard this TP metric as a more reliable indicator of genuine mathematical problem-solving ability. For high-performing models such as o4-mini, DeepSeek-R1, and Gemini2.5-Pro-Preview, TP scores in the reasoning & arithmetic confusion matrix closely match their ACC@Fi values in Tab. 1. In contrast, GPT-4.1 and Claude3.7-Sonnet exhibit significantly lower TP scores relative to their ACC@Fi, suggesting that their final answer accuracy may overestimate their true reasoning capabilities.
5 Conclusion
We introduce the SMART evaluation framework and its accompanying benchmark, SMART-Bench, for systematic assessment of LLMs in mathematical problem solving. SMART decomposes this complex ability into four interpretable dimensions, Understanding, Reasoning, Arithmetic, and Reflection & Refinement, enabling more granular evaluation than traditional answer-based metrics. A key feature is its self-generating and self-validating mechanism, which iteratively verifies the correctness of generated test data to ensure reliability and scalability. Experiments on 21 open- and closed-source LLMs reveal that reasoning and reflection remain primary bottlenecks. Moreover, our analysis highlights the limitations of final-answer accuracy as a sole indicator and introduces a new metric to better capture true model capability. We hope SMART serves as a foundation for more comprehensive and interpretable evaluations.
6 Limitation
While our proposed framework provides a comprehensive evaluation platform, it is important to acknowledge its inherent limitations. In particular, although Z3 and SMT-LIB effectively handle linear, integer, and some nonlinear constraints, their problem-solving capabilities are restricted. They are unsuitable for highly complex nonlinear problems and certain NP-complete combinatorial tasks. To overcome these limitations, future work will investigate using Lean Moura and Ullrich (2021) to formalize and prove complex mathematical theorems involving higher-order logic and intricate proof structures beyond SMT solvers’ scope.
Appendix A Polya’s Problem-solving Theory
Polya first propose the four step problem solving method in the book ’How to solve it’ Pólya and Conway (1957) to systematically answer questions. According to his theory, it breaks down the process of problem-solving into four steps (Fig.  1): 1) understanding the problem. 2) make a plan. 3) carry out the plan. 4) look back and review.
Inspired by Polya’s problem-solving theory, we evaluate the mathematical capability in four key dimensions:
  • The understanding dimension assesses the model’s ability to comprehend the problem accurately and clearly.
  • The reasoning dimension evaluates the model’s capacity for symbolic reasoning and logical deduction.
  • The arithmetic dimension measures the model’s proficiency in performing numerical computations.
  • The reflection & refinement dimension examines the model’s ability to identify errors in the solution process and make corrections to improve the final answer.
Appendix B SMART-Bench Generation
B.1 Seed Question Dataset of SMART
The foundation of the SMART benchmark is a seed dataset comprising 2000 problem instances. These were curated from seven widely-used mathematical problem datasets: GSM8K Cobbe et al. (2021), SVAMP Patel et al. (2021), ASDiv Miao et al. (2020), AQuA Ling et al. (2017), MAWPS Koncel-Kedziorski et al. (2016), MATH Hendrycks et al. (2021), and problems from the AIME 2024 competition Huggingface (2024). These initial 2,000 seed questions, along with their subsequently generated dimension-specific variations (four per seed question), form the complete SMART-Bench, totaling 10,000 test instances.
Several criteria were applied during the selection and processing of these seed questions. To ensure consistency in question format, problems from the AQuA dataset, originally multiple-choice, were converted into an open-ended format; the textual content of the correct option was adopted as the ground truth for these transformed questions. Furthermore, we excluded problems whose solutions fundamentally rely on operations not readily expressible or automatically verifiable using SMT-LIB, such as calculations involving the greatest common divisor (GCD), least common multiple (LCM), or the determination of maximum/minimum values within a set. Questions requiring multiple distinct numerical values in their answers were also omitted. To maintain a baseline level of difficulty and focus on multi-step problem-solving, we filtered out mathematical problems that could be solved in a single reasoning step. The distribution of reasoning steps for the selected 2,000 seed questions is depicted in Fig. 5, which illustrates that the majority of problems in SMART-Bench involve multiple reasoning steps, with a notable concentration in the 2 to 7 step range.
Fig.  6 presents a data sample in the SMART benchmark, which contains the seed question, the extracted context, the SMT-LIB expression, the arithmetic notation question, the CoT and the final answer.
B.2 Dimension-Specific Testing Data and Task Description
For each seed question, SMART generates distinct variants to evaluate the four targeted problem-solving dimensions. The generation and ground truth creation for each dimensional task are described below.
B.2.1 Understanding
To generate ground truth for the understanding dimension, we utilize GPT-4o to perform context extraction. The extracted context is structured into the following components: Problem Scenario (describing the overall context of the problem), Goal (specifying what needs to be solved), Known Quantities (listing explicitly provided numerical values or facts), Unknown Quantities (identifying variables or values to be determined), Relationships and Constraints (detailing connections and limitations between pieces of information), and Irrelevant Information (pinpointing details not pertinent to the solution). The prompt employed to guide GPT-4o in extracting this contextual information for ground truth generation is presented in Fig. 14. To enhance the performance of GPT-4o in generating these contexts, we employed a three-shot prompting strategy. To ensure the quality of this generated ground truth, we adopt the human verification and iterative regeneration strategy outlined by Chen et al. (2024). Specifically, after the initial generation of ground truth contexts, a random 10% sample of these is manually assessed by human annotators. If the annotators find less than 80% of the sampled contexts satisfactory, the entire set of contexts is regenerated using a refined approach, and the verification process is repeated.
B.2.2 Reasoning
For the Reasoning dimension, the task does not involve generating new questions or altering the ground-truth answer from the seed question. Instead, LLMs are tasked with generating a symbolic representation of the original seed question, typically in SMT-LIB format. The correctness of this generated symbolic formula is then indirectly assessed: the formula is executed by an SMT solver, and the result obtained from the solver is compared against the known ground-truth answer of the original seed question. A match indicates a correct reasoning process as captured by the SMT-LIB formula.
B.2.3 Arithmetic
The arithmetic ability of LLMs is measured by their performance on pure numeric calculation problems that share the same reasoning logic and final answer as the seed question. Directly converting a seed question into an arithmetic notation problem is challenging for LLMs. This difficulty arises because it requires simplifying complex natural language into structured mathematical operations while preserving the logical relationships between variables. Such a transformation is non-trivial, as the model must accurately interpret the problem’s intent, handle ambiguous phrasing, and correctly map linguistic constructs to arithmetic operations. Therefore, to address this, we first generate an SMT-LIB representation of the seed question, which simplifies the reasoning logic among variables. Subsequently, we utilize GPT-4o to convert this SMT-LIB representation into the arithmetic notation problem. The prompt for that process is shown in Fig.  7 and Fig.  8.
B.2.4 Reflection & Refinement
For the Reflection & Refinement dimension, we first generate Chain-of-Thought (CoT) solutions containing deliberate errors. To create these erroneous CoTs, one of the six error types defined in Fig. 18 (e.g., arithmetic number error, skipped step) is uniformly sampled and injected into a correct CoT. The ground truth for the reflection sub-task is then the specific name of the introduced error type and its order number. For the subsequent refinement sub-task, direct verification of the LLM-corrected CoT is complex. Instead, we evaluate the refined CoT by using GPT-4o to generate a final answer based upon it. The LLM is considered to have successfully passed the refinement task if this GPT-4o-generated answer matches the ground truth of the original seed question.
Figure 5: The number of reasoning step statistics of seed question dataset.
Figure 6: A data sample in the SMART benchmark.
Appendix C SMART-Bench Verification
We adopt the neuro-symbolic system to automatically verify the correctness of generated SMT-LIB and its arithmetic notation formed question. This self-validating mechanism can also be use to verify the augmented version of questions which are generated based on the number of operator modification rewriting rules. The process is as following:
  1. 1.
    Initial SMT-LIB Conversion and Validation: We first attempt to convert the original seed question  into an SMT-LIB formula. We employ an SMT solver to compute the solution from this SMT-LIB formula. If the solver-derived result does not match the known ground-truth answer of the original seed problem, the SMT-LIB generation process is iteratively refined and re-verified until consistency is achieved. This ensures a validated symbolic representation of the original problem.
  2. 2.
    Augmentation in SMT-LIB Space: Once a validated SMT-LIB formula for the original problem  is obtained, numerical or operational changes dictated by our rewriting rules are applied directly within this symbolic SMT-LIB space.
  3. 3.
    Automated Ground-Truth Generation for Augmented Problems: The modified SMT-LIB formula, now representing an augmented version of the problem, is then solved using the Z3-solver. The result provided by the solver serves as the automatically generated and verified ground-truth answer for this new, dynamically created problem instance.
Figure 7: The prompt for LLMs to convert the seed question to a symbolic expression.
Figure 8: The prompt for LLMs to convert the SMT-LIB expression to arithmetic notation problem.
Figure 9: The prompt for LLMs to solve the arithmetic notation problem.
Figure 10: The prompt for LLMs to detect mistakes in the CoT.
Figure 11: The prompt for LLMs to detect mroe than one mistakes in the CoT.
Figure 12: The prompt for LLMs to correct the mistaks in the CoT.
Figure 13: The prompt for LLMs to refine the anwer of the seed question based on the correct CoT.
Appendix D Experiments
D.1 Experiment Setting
All experiments were conducted on a Linux server equipped with two NVIDIA H800 GPUs (80GB). The GPUs were used for deploying and performing inference on open-source models. The Python version used was 3.9.20, and the version of the Transformers package was 4.46.0.
D.2 Prompt and Evaluation metrics in SMART
D.2.1 Understanding
We propose a context extraction task to measure the capability of LLMs in understanding math problems. The input for this task consists of seed questions, and the model is asked to extract the context of the question based on the prompt shown in Fig. 14.
We employed an LLM-as-a-judge method Zheng et al. (2023) to evaluate the quality of generated context. In this method, GPT-4o is used impartially to assess semantic similarity between the background context extracted by the evaluated LLM and the reference ground-truth, assigning a score from 1 to 100.
Figure 14: The prompt for LLMs to extract context from seed question.
D.2.2 Reasoning
For the reasoning dimension, we introduce a symbolic formalization task to evaluate the symbolic reasoning capability of LLMs. The input for this task is the seed question, and LLMs are asked to generate the SMT-LIB expression of the question, without solving the problem. The prompt for this task is shown in Fig. 7. Subsequently, the Z3 solver is used to compute the results of the generated SMT-LIB expression. The evaluation metric for this task is the accuracy of the SMT-LIB formula results. Specifically, the accuracy of generating correct SMT-LIB expressions, determined by the number of correct SMT-LIB (verified by Z3-solver) divided by the total number of problems.
D.2.3 Arithmetic
For the arithmetic dimension, we introduce a numeric calculation task to assess the arithmetic ability of LLMs. The input for this task is an arithmetic notation problem, and LLMs are asked to solve it using the prompt shown in Fig. 9. The arithmetic problem is designed to have the same answer and reasoning logic as its corresponding seed question. The evaluation metrics for this task is accuracy, defined as the number of arithmetic problems solved correctly divided by the total number of problems.
D.2.4 Reflection & Refinement
For the reflection & refinement dimension, we propose an error correction task that requires LLMs to detect mistakes in the chain-of-thought (CoT) of seed questions, correct these mistakes, and generate a new answer for the seed question. This task consists of three steps. The first step involves detecting errors in the CoT, given the question and CoT, with the answer being the the specific name of the introduced error type and its order number, as specified in the prompt shown in Fig. 10 and prompt for more errors in Fig. 11. If LLMs fail to detect all mistakes, they do not need to attend the following refinement task. The second step is to fix errors in CoT and generate corrected CoT with given seed question and CoT with prompt shown in Fig.12. The final step is to refine the answer based on the corrected CoT and the seed question, using the prompt in Fig. 13. If LLMs successfully detect all mistakes and generate the correct final answer based on the corrected CoT, we consider the model to have passed the error correction task.
The evaluation metrics for the reflection task is accuracy. The accuracy in identifying all errors in the chain-of-thought (CoT) solution, computed as the number of problems where all CoT errors are correctly identified divided by the total number of problems. The accuracy of refining the CoT after error identification, measured as the number of problems correctly answered based on the refined CoT divided by the number of problems with correctly identified CoT errors. The combined accuracy for both reflection and refinement tasks, calculated as the number of problems correctly answered after CoT refinement divided by the total number of problems.
D.3 Difficulty setting for Dimension-specific Task
To create dimension-specific questions with varying difficulty levels, we employ the following strategies:
For the understanding dimension, difficulty is controlled by progressively introducing irrelevant sentences, sourced from other problems, as ’noise’ within the seed question’s text. The number of such noise sentences dictates the complexity of the context extraction task. The ground truth for these modified questions is updated by incorporating these noise sentences into the ’Irrelevant Information’ category of the context extraction template.
For the reasoning dimension, question complexity is defined by the number of distinct mathematical operations (e.g., +,,×,÷$+,-,\times,\div,mod$) required to formulate the solution. Problems are then categorized into multiple difficulty levels based on this operational count.
In the arithmetic dimension, complexity is varied by altering the number of digits in the numerical values involved (e.g., changing ’12’ to a five-digit number like ’34.823’), rather than solely their magnitude, as precision with more digits presents a distinct challenge. The ground truth for these modified arithmetic problems is obtained through our self-verification mechanism.
In the reflection & refinement dimension, difficulty is modulated by the number and types of mistakes deliberately injected into the Chain-of-Thought solutions. We randomly introduce a varying number of distinct error types (from the categories listed in Fig. 18) into the CoT to create different levels of challenge for the error detection and correction tasks.
Each seed question undergoes a two-stage transformation process. In the first stage, it is decomposed into four distinct, dimension-specific tasks that enable fine-grained evaluation of individual capabilities. In the second stage, these tasks are further rewritten using validated augmentation strategies to generate diverse variants that test the robustness and adaptability of LLMs. This hierarchical, two-phase generation framework enhances reliability and scalability by enabling comprehensive and granular evaluation while mitigating risks of overfitting and data contamination.
D.4 Examples for question with different difficulty setting
Fig.15 shows examples of different difficulty settings for the understanding dimension evaluation. The sentences with a red background in the image represent irrelevant noise sentences, and the more noise sentences there are, the harder the task of extracting the effective context becomes.
Fig.16 shows examples of question with different reasoning steps, which indicates different reasoning difficulties.
Fig.17 presents arithmetic questions with numbers in different digits. Numbers with more digits means more difficult for arithmetic evaluation task.
Fig.18 presents CoT with different type errors. CoT with more mistakes are more difficult for reflection and refinement evaluation task.
Figure 15: Example of questions with different number of noise sentences.
Figure 16: Example of questions with different reasoning steps.
Figure 17: Example of arithmetic questions with number in different digits.
Figure 18: Example of CoT with different errors.
D.5 Zero-shot vs. three-shot
We conducted additional comparative experiments evaluating model performance with and without three-shot examples across all stages. The results are presented in the tab.3. As shown in the table, removing the three-shot examples leads to a significant performance drop across all models for dimensions. This confirms the substantial positive impact of including few-shot examples for these more complex reasoning stages. Our comparative results show that adding few-shot examples provides an improvement in all dimension scores across.
Model Final Answer Understanding Reasoning Arithmetic Reflection&Refinement
Zero-shot Three-shot Zero-shot Three-shot Zero-shot Three-shot Zero-shot Three-shot Zero-shot Three-shot
GPT-4.1 43.25 59.34 90.25 95.41 59.52 73.29 38.65 45.71 37.05 59.90
o4-mini 92.25 93.75 93.58 95.22 85.63 90.72 97.12 98.25 51.20 55.05
Gemini2.5-Flash-Preview 90.65 93.33 91.12 94.25 81.47 89.72 95.41 98.18 30.74 35.85
Table 3: Zero-shot and three-shot prompt engineer strategy comparison in SMART.

### v3

@@
 Abstract
-Large Language Models have achieved remarkable results on a variety of mathematical benchmarks.
-However, concerns remain as to whether these successes reflect genuine mathematical reasoning or superficial pattern recognition.
-Common evaluation metrics, such as final answer accuracy, fail to disentangle the underlying competencies involved, offering limited diagnostic value.
-To address these limitations, we introduce SMART: a Self-Generating and Self-Validating Multi-Dimensional Assessment Framework.
-SMART decomposes mathematical problem solving into four distinct dimensions: understanding, reasoning, arithmetic, and reflection & refinement.
+Large Language Models (LLMs) have achieved remarkable results on a variety of mathematical benchmarks.
+However, concerns remain as to whether these successes reflect genuine reasoning or superficial pattern recognition.
+Common evaluation methods, which focus on the either the final answer or the reasoning process, fail to assess the entire problem-solving procedure.
+To address these limitations, we introduce SMART: a Self-Generating and Self-Validating Multi-Dimensional Assessment Framework, together with its corresponding benchmark, SMART-Bench.
+SMART decomposes the entire problem solving process into four distinct cognitive dimensions: Understanding, Reasoning, Arithmetic, and Reflection & Refinement.
 Each dimension is evaluated independently through tailored tasks, enabling interpretable and fine-grained analysis of LLM behavior.
-Crucially, SMART integrates an automated self-generating and self-validating mechanism to produce and verify benchmark data, ensuring both scalability and reliability.
 We apply SMART to 21 state-of-the-art open- and closed-source LLMs, uncovering significant discrepancies in their abilities across different dimensions.
-Our findings demonstrate the inadequacy of final answer accuracy as a sole metric and motivate a new holistic metric to better capture true problem-solving capabilities.
+Our findings reveal genuine weaknesses in current LLMs and motivate a new metric, the All-Pass Score, to better capture true problem-solving capabilities.
 Code and benchmarks will be released upon acceptance.
-SMART: Self-Generating and Self-Validating Multi-Dimensional Assessment for LLMs’ Mathematical Problem Solving
-Yujie Hou, Ting Zhang, Mei Wang, Xuetao Ma, Hua Huangthanks:  Corresponding author: Hua Huang School of Artificial Intelligence, Beijing Normal University, Beijing, China {houyujie, maxuetao}@mail.bnu.edu.cn, {tingzhang,wangmei1, huahuang}@bnu.edu.cn
-1 Introduction
-Large language models (LLMs)  Achiam et al.
-(2023); Wei et al.
-(2022) have demonstrated impressive performance across various natural language processing tasks Singhal et al.
-(2023); Wang et al.
-(2024); Jiang et al.
-(2024); Ma et al.
-(2025) and are being increasingly integrated into real-world applications, such as education, scientific computing, and decision support OpenAI (2024); Guo et al.
-(2025).
-As their adoption accelerates, it is critical to rigorously evaluate their true capabilities.
-Figure 1: Following Polya’s four-step framework for human mathematical problem solving—understanding the problem, devising a plan, carrying out the plan, and looking back—we define four corresponding evaluation dimensions and design specific tasks to systematically assess the capabilities of large language models.
-While numerous benchmarks have been proposed to evaluate the mathematical reasoning capabilities of LLMs  Cobbe et al.
-(2021); Ling et al.
-(2017); Patel et al.
-(2021); Miao et al.
-(2020); Koncel-Kedziorski et al.
-(2016), many suffer from a significant limitation: they provide insufficient interpretability regarding actual mathematical proficiency.
-The predominant focus on final answer accuracy is particularly problematic, as this metric reflects an aggregate outcome influenced by multiple, intertwined cognitive processes.
-As a result, existing benchmarks fail to effectively evaluate the distinct components of the mathematical problem-solving process Cobbe et al.
-(2021); Ling et al.
-(2017); Patel et al.
-(2021); Miao et al.
-(2020); Koncel-Kedziorski et al.
-(2016).
+Introduction
+Large language models (LLMs)  (Achiam et al.
+2023; Wei et al.
+2022) have demonstrated impressive performance across various natural language processing tasks (Singhal et al.
+2023; Wang et al.
+2024; Jiang et al.
+2024; Ma, Jiang, and Huang 2025) and are being increasingly integrated into real-world applications, such as education, scientific computing, and decision support (OpenAI 2024; Guo et al.
+2025).
+With the increasing adoption of LLMs, it is critical to rigorously evaluate their true capabilities.
+Figure 1: Comparison between SMART and existing methods.
+Final answer-based evaluations assess only the correctness of the final outcome.
+Reasoning process-based evaluations focus on identifying errors within the solution path.
+Our proposed SMART method decomposes the entire problem-solving process into four distinct dimensions, each evaluated through targeted tasks.
+This enables a more fine-grained and systematic assessment of LLMs’ mathematical problem-solving abilities.
+Mathematical questions are widely used as proxy tasks to evaluate the cognitive abilities of LLMs.
+Although numerous mathematical benchmarks have been proposed, several important limitations persist.
+A primary issue is that current datasets do not holistically assess the entire problem-solving process, which involves multiple, intertwined cognitive dimensions.
+For example, widely used benchmarks such as GSM8K (Cobbe et al.
+2021) and MATH (Hendrycks et al.
+2021) focus primarily on final answer accuracy, while others—such as MR-GSM8K (Zeng et al.
+2024a)—target specific aspects like the reasoning process.
+A comprehensive process-level evaluation is essential for measuring the true capabilities of LLMs in mathematical problem solving.
 Another critical deficiency stems from the static nature of most existing benchmarks.
-Such static evaluation sets are highly susceptible to data contamination, potentially leading to inflated performance scores and skewed assessments of LLM capabilities Oren et al.
-(2023); Zhu et al.
-(2023, 2024); Li et al.
-(2024a).
+Such static evaluation sets are highly susceptible to data contamination, potentially leading to inflated performance scores and skewed assessments of LLM capabilities (Oren et al.
+2023; Zhu et al.
+2023, 2024; Li et al.
+2024a).
 Furthermore, as LLMs rapidly advance, they can quickly reach performance saturation on these fixed datasets.
-Consequently, static benchmarks may no longer accurately reflect the true capabilities of more advanced models Fodor (2025); AI@Meta (2024).
-Furthermore, constructing high-quality benchmark datasets is resource-intensive, requiring significant human labor and time for annotation Mirzadeh et al.
-(2024); Kurtic et al.
-(2024).
-When leveraging advanced LLMs like GPT-4o for data generationLi et al.
-(2024b); Zhu et al.
-(2024), ensuring the accuracy and consistency of the generated content remains a major challenge, as LLMs may produce errors and inconsistencies.
-Although human annotators are often employed to review and refine LLM-generated questions and answers Zhu et al.
-(2024); Zheng et al.
-(2023), this manual intervention is costly and inherently limits scalability.
+As a result, static benchmarks may no longer accurately reflect the true capabilities of more advanced models (Fodor 2025; AI@Meta 2024).
+Furthermore, constructing high-quality mathematical benchmarks demands considerable human effort and annotation resources.
+For instance, miniF2F (Zheng, Han, and Polu 2021) and FIMO (Liu et al.
+2023) manually formalize IMO-level problems with carefully curated formal statements or solutions.
+Similarly, MR-BEN (Zeng et al.
+2024b) involves costly step-level annotations.
+As the number of samples increases, the annotation cost rises dramatically, significantly limiting the scalability of such benchmarks.
 Figure 2: An overview of the SMART framework for evaluating the mathematical problem-solving process.
@@
 Each dimension is evaluated using dimension-specific task and metrics, ensuring a comprehensive assessment of the model’s problem-solving capabilities.
-In light of these issues, we propose a novel evaluation framework: the Self-Generating and Self-Validating Multi-Dimensional Assessment Framework (SMART).
-Inspired by Polya’s problem-solving theory Polya (2014), SMART deconstructs the mathematical problem-solving process into four key evaluation dimensions: understanding, reasoning, arithmetic, and reflection & refinement.
-Each dimension is designed to evaluate a distinct facet of problem-solving progression (Fig.
+The yellow-highlighted symbolic assertion illustrates an inferred condition that is not explicitly stated in the original problem.
+In light of these limitations, we propose a novel evaluation framework: the Self-Generating and Self-Validating Multi-Dimensional Assessment Framework (SMART), along with a new metric called the All-Pass Score.
+Inspired by Polya’s problem-solving theory (Polya 2014), SMART decomposes the mathematical problem-solving process into four key evaluation dimensions: Understanding, Reasoning, Arithmetic, and Reflection & Refinement (Fig.
 1).
-To implement this multi-dimensional assessment, SMART generates dimension-specific tasks with clearly defined outputs from seed problems.
-This structured decomposition allows SMART to enable the model to probe distinct cognitive components of mathematical problem solving.
-By isolating each dimension, SMART facilitates comprehensive evaluation and yields insights into the strengths and weaknesses of LLMs, informing targeted algorithmic enhancements.
-The resulting evaluation dataset, SMART-Bench, is built through an integrated self-generating and self-validating mechanism that not only programmatically produces dimension-specific data but also iteratively verifies its quality.
-This approach ensures both scalability and reliability in benchmark construction.
+This evaluation paradigm enables a comprehensive assessment of the entire problem-solving process—from initial problem comprehension, to planning reasoning steps, performing numerical computations, and finally checking the solution.
+To implement this multi-dimensional assessment, we introduce a self-generating and self-validating mechanism that dynamically constructs dimension-specific tasks from seed problems, forming our benchmark SMART-Bench.
+These dimension-specific tasks are novel to LLMs and thus help alleviate data contamination.
+We adopt a neuro-symbolic framework (Barrett et al.
+2010; De Moura and Bjørner 2008) to iteratively verify the correctness of the generated questions, significantly reducing the need for manual annotation.
+Thus, our approach ensures the scalability and reliability of benchmark construction.
 We evaluate 21 recently released open- and closed-source LLMs on SMART-Bench.
-Our experimental results reveal that models exhibit diverse performance patterns, demonstrating varied strengths and weaknesses across different dimensions.
-Furthermore, our investigation pinpoints key factors influencing performance across these distinct dimensions.
-Finally, through an in-depth analysis of the relationship between final answer accuracy and the scores of our decomposed dimensions, we propose a new metric designed to more accurately measure the true mathematical capabilities of LLMs.
+Experimental results demonstrate that even the most advanced models perform poorly under the proposed All-Pass Score, underscoring the challenging nature of our benchmark.
+Notably, the Reasoning and Reflection dimensions emerge as the primary bottlenecks in mathematical problem solving.
+Additionally, SMART serves as a diagnostic tool, identifying which dimensions are responsible for performance drops under dynamic evaluation.
+Furthermore, we find that targeted improvements in specific weak dimensions can lead to substantial overall gains in mathematical capability—for example, increasing the final answer accuracy of Qwen2.5-72B by 11.77%.
 Our main contributions are as follows:
   1. $\bullet$
-    We propose SMART, a self-generating and self-validating multi-dimensional assessment framework designed to break down the mathematical problem-solving process of LLMs into four key dimensions.
-This approach enables deeper insight into underlying cognitive processes beyond final answer accuracy.
+    We propose the SMART evaluation paradigm along with a novel All-Pass Score metric to enable fine-grained and comprehensive assessment of a model’s mathematical problem-solving abilities.
   2. $\bullet$
-    We introduce an integrated self-generating and self-validating mechanism that produces SMART-bench dataset.
-This significantly reduces reliance on human annotation while ensuring the benchmark’s scalability and reliability.
+    We introduce an integrated self-generating and self-validating mechanism for constructing SMART-Bench.
+The self-generating component creates novel tasks from seed problems to mitigate data contamination, while the self-validating component verifies correctness to reduce human annotation.
+Together, they ensure the scalability and reliability of the benchmark.
   3. $\bullet$
@@
 These findings offer a valuable foundation for future LLM development and optimization.
-2 Related Work
-Mathematic benchmark.
-Numerous mathematical benchmarks with varying levels of difficulty have been developed to explore the upper bound of LLMs’ mathematical capabilities.
-These benchmarks range from grade-school-level datasets Cobbe et al.
-(2021), to high-school-level datasets Hendrycks et al.
-(2021), and extend to expert-level datasets Glazer et al.
-(2024).
+Related Work
+Mathematic benchmark Numerous mathematical benchmarks with varying levels of difficulty have been developed to explore the upper bound of LLMs’ mathematical capabilities.
+These benchmarks range from grade-school-level datasets (Cobbe et al.
+2021), to high-school-level datasets (Hendrycks et al.
+2021), and extend to expert-level datasets (Glazer et al.
+2024).
 Their scope covers a broad range of mathematical domains, including geometry, number theory, and real analysis.
-However, despite the increasing difficulty of these benchmarks, it remains unclear whether LLMs truly comprehend mathematical concepts or merely rely on pattern-matching to generate correct answers Mirzadeh et al.
-(2024).
-Furthermore, the widespread use of these benchmarks increase the risk of data contamination, potentially inflating performance evaluations Li et al.
-(2024a).
-Dynamic evaluation.
-To address these concerns, recent studies have explored dynamic evaluation approaches Zhu et al.
-(2023, 2024), which generate adaptive test data based on pre-defined rewriting rules.
-For example, GSM-Plus Li et al.
-(2024b) introduces eight perturbation strategies, while GSM-Symbolic Mirzadeh et al.
-(2024) generates diverse problem variants using symbolic templates.
-These approaches aim to mitigate data leakage and improve robustness in assessments.
-Despite these advancements, current evaluation methodologies remain limited.
-They predominantly focus on final-answer accuracy, which fails to capture an LLM’s reasoning process in a systematic manner.
-Additionally, they do not explicitly assess the diverse cognitive skills required for mathematical problem-solving, nor do they provide fine-grained insights into the underlying logical reasoning steps.
-These limitations underscore the need for a comprehensive, interpretable, and scalable evaluation framework that can decompose the problem-solving process, assess reasoning at multiple levels, and minimize reliance on human verification.
-To bridge this gap, we introduce SMART, a self-validating multi-dimensional assessment framework designed to systematically evaluate LLMs’ mathematical reasoning abilities while enhancing benchmark reliability and scalability.
-3 Methodology
-3.1 SMART Framework
-Understanding.
-The initial step in mathematical problem-solving is to comprehend the problem.
-Accordingly, the first dimension in the SMART framework assesses the ability of LLMs to clearly and comprehensively understand a given problem before attempting a solution.
-To rigorously evaluate this dimension, we introduce a context extraction task.
-This task requires the LLM to meticulously identify and categorize various informational components embedded within the problem.
+However, despite their increasing difficulty, these benchmarks primarily adopt an outcome-based evaluation approach, making it unclear whether LLMs genuinely understand mathematical concepts or simply rely on pattern-matching to produce correct answers (Mirzadeh et al.
+2024).
+To address this, ProcessBench (Zheng et al.
+2025) and PRMBench (Song et al.
+2025) have been proposed to enable process-level evaluation by identifying erroneous steps in the model’s mathematical reasoning.
+Nevertheless, these process-level benchmarks still fall short of evaluating the entire spectrum of human mathematical thinking, as they do not comprehensively assess all stages of the problem-solving process
+Dynamic evaluation The widespread use of benchmarks increase the risk of data contamination, potentially inflating performance evaluations (Li et al.
+2024a).
+To address these concerns, recent studies have explored dynamic evaluation approaches (Zhu et al.
+2023, 2024), which generate adaptive test data based on pre-defined rewriting rules.
+GSM-Plus (Li et al.
+2024b) and GSM-Symbolic (Mirzadeh et al.
+2024) generates problem variants from seed questions.
+These approaches aim to mitigate data leakage and improve robustness in evaluations.
+However, manually annotating the newly generated questions remains highly labor-intensive and costly.
+This motivates the need for automated mechanisms that can generate and verify evaluation data at scale.
+These limitations underscore the need for a framework that can comprehensively assess the entire problem-solving process, provide fine-grained and interpretable analyses, and reduce the cost of benchmark construction.
+To address this gap, the SMART framework is designed to systematically evaluate the mathematical reasoning abilities of LLMs.
+Dimension Input question Input source Output Ground truth Ground truth source Task evaluator
+Understanding Seed question / Structured information Structured information
+GPT-4.1
++ human verification
+LLM-as-a-judge
+Reasoning Seed question / SMT-LIB
+Final answer of
+seed question
+Z3 Solver and
+Rule-match
+Arithmetic
+Arithmetic
+notation question
+SMT-LIB+GPT-4.1
++ human verification
+Answer of
+arithmetic question
+Final answer of
+seed question
+/ Rule-match
+Reflection
+CoT solution with
+injected errors
+Rule Error types Error types / Rule-match
+Refinement
+CoT solution with
+injected errors
+Rule Corrected CoT
+Final answer of
+seed question
+/ Rule-match
+Table 1: Overview of Dimension-specific Task Design and Ground truth in SMART
+Methodology
+Evaluation Dimensions
+Understanding
+Understanding the problem is the first step in mathematical problem solving.
+Accordingly, the first dimension of SMART evaluates an LLM’s understanding ability through an information extraction task, which requires the model to accurately identify and categorize key components within the problem statement.
 As illustrated in Fig.
-2 (Dimension 1, Task 1), the LLM is prompted to extract and structure key information into a predefined template.
-This template includes categories such as: problem scenario, goal, known and unknown quantities, relationships and constraints, and irrelevant information.
+2 Dimension 1, the LLM is prompted to extract and organize this information into a predefined template.
+The template includes categories such as the problem scenario, goal, known and unknown quantities, relationships and constraints, and irrelevant information.
 This structured context extraction task serves a dual purpose.
 At a surface level, it evaluates the LLM’s ability to summarize and extract salient information.
-More profoundly, it probes the depth of the LLM’s comprehension by requiring it to discern the roles and interconnections of different problem elements, thereby offering a nuanced measure of its problem understanding capabilities.
-This moves beyond simple keyword spotting to assess a more holistic grasp of the problem’s structure.
-Reasoning.
-Subsequent to the comprehension phase, Polya’s problem-solving heuristic emphasizes devising a coherent plan.
-Aligning with this, the second dimension evaluates the reasoning ability of LLMs.
-This is achieved through a symbolic formalization task, wherein LLMs are required to solve mathematical problems via formal symbolic expressions.
-This approach compels the LLM to concentrate on the inherent logical structure of the problem and the intricate relationships between its constituent elements, rather than relying on direct numerical computation.
-Beyond merely translating problem statements into symbolic form, this task demands the identification of solution-relevant equations and the construction of a logically consistent path to the answer.
+More profoundly, it probes the depth of the LLM’s comprehension by requiring it to discern the roles and interconnections of different elements, thereby offering a nuanced measure of its problem understanding capabilities.
+Reasoning
+Following the comprehension phase, the next stage involves devising a reasoning plan.
+Thus, the second dimension of SMART evaluates the reasoning ability of LLMs through a symbolic formalization task, in which models are required to solve mathematical problems using formal symbolic expressions in SMT-LIB format (Barrett et al.
+2010).
+This approach compels the model to focus on the underlying logical structure of the problem and the intricate relationships among its components, rather than relying on direct numerical computation.
+Although the SMT-LIB format is seldom used in LLMs, experiments presented in the Appendix demonstrate that the primary challenge of the Reasoning task lies in generating a correct reasoning path, rather than translating natural language into formal language.
+The reasoning task extends beyond merely translating problem statements into symbolic form; it requires LLMs to identify solution-relevant equations and construct a logically coherent path to the answer.
 As illustrated in Fig.
-2 (Dimension 2, Task 2), a highlighted symbolic assertion, absent from the original question, captures a key inferred relation, indicating the model’s ability to derive implicit knowledge.
-This dimension thus probes deeper reasoning by emphasizing inference and the construction of a complete logical model.
-We adopt a standardized and widely used format for representing satisfiability modulo theories (SMT) problems Barrett et al.
-(2010).
-The SMT-LIB expressions produced by the models are subsequently verified using established SMT solvers such as Z3 De Moura and Bjørner (2008), which can automatically determine the satisfiability of logical formulas or compute concrete solutions.
-A reasoning plan expressed in SMT-LIB is considered correct if the solution derived by the solver consistently align with the ground truth of the original mathematical problem.
-Arithmetic.
+2 (Dimension 2), the highlighted symbolic assertion is not explicitly provided as a known condition in the original question.
+Instead, LLMs must infer this key constraint to solve the problem, thereby demonstrating their ability to derive implicit knowledge.
+Arithmetic
 The third dimension, Arithmetic, is dedicated to assessing an LLM’s proficiency in performing accurate numerical computations.
-This evaluation is conducted through a numeric calculation task that presents problems exclusively in terms of numerical values and variables, requiring the execution of fundamental arithmetic operations.
-For instance, as illustrated in Fig.
-2 (Dimension 3, Task 3), the arithmetic questions are designed to contain only basic calculation operations and clearly defined variables.
-This task design isolates arithmetic skills by minimizing confounding cognitive demands such as language comprehension or complex reasoning, thereby enabling a focused and accurate evaluation of the model’s computational proficiency.
-Reflection & Refinement.
-The final stage in Polya’s problem-solving model, "looking back," involves reviewing the solution and improving its correctness.
-Recent literature has demonstrated that LLMs possess a promising capacity for self-feedback, and iterative refinement Madaan et al.
-(2023); Shinn et al.
-(2023); Guo et al.
-(2025); Huang et al.
-(2022).
-These studies suggest that prompting LLMs to scrutinize their own outputs or follow structured verification steps lead to improved accuracy and more robust reasoning.
-So, it is crucial to evaluate the self-critique capability of LLMs.
-Accordingly, the fourth dimension of our SMART framework, Reflection & Refinement, aims to evaluate the LLM’s ability to critically review its problem-solving process and refine its answers.
-This is operationalized through an error correction task, an aspect largely overlooked in prior benchmarks.
-As depicted in Fig.
-2 (Dimension 4, Task 4), LLMs are presented with a Chain-of-Thought (CoT) solution that contains deliberately introduced errors.
-We introduce six distinct categories of errors into the CoT, encompassing arithmetic number error, skipped step error, hallucinatory injection error, logical order error, redundant output error, and incorrect operator error (detailed in Fig.
-18).
-The LLM is then tasked with first identifying any flaws within the provided CoT, then correcting these identified errors, and finally generating a refined answer based on the corrected reasoning.
-A critical design choice is that if an LLM fails to detect any errors in the provided CoT, it does not proceed to the subsequent refinement tasks for that particular problem instance.
-3.2 The Self-Generating and Self-Validating Mechanism in SMART
-A core innovation of SMART lies in its integrated self-generating and self-validating mechanism, which ensures both scalable benchmark construction and the reliability of SMART-Bench data.
-3.2.1 Dataset Generation
-To construct SMART-Bench, we begin by curating a diverse set of 2,000 seed questions, selected from seven widely used mathematical problem datasets: GSM8K Cobbe et al.
-(2021), SVAMP Patel et al.
-(2021), ASDiv Miao et al.
-(2020), AQuA Ling et al.
-(2017), MAWPS Koncel-Kedziorski et al.
-(2016), MATH Hendrycks et al.
-(2021), and the AIME 2024 competition Huggingface (2024).
-For each seed question, we automatically generate four dimension-specific variants, each aligned with a core dimension.
-The resulting dataset comprises 10,000 instances, 2,000 original questions and 8,000 generated variants.
-Details about generation procedures are provided in Appendix B.
-3.2.2 Dataset Verification
-For the understanding dimension, given the strong capability of LLMs in information extraction, we rely on a robust LLM to ensure correctness.
-In contrast, for reasoning and arithmetic questions, manual verification by human annotators is resource-intensive and limits scalability.
-To overcome this, SMART incorporates a self-validating mechanism based on a neuro-symbolic framework that automatically and precisely verifies generated question variants and their answers for dimensions suitable for symbolic representation.
-For example, to verify the correctness of a generated SMT-LIB expression for a seed question, we employ an SMT solver to compute the symbolic formula’s answer automatically.
-We then compare the solver’s output with the known seed answer, iterating this process until the generated SMT-LIB expression aligns with the expected result.
-Further details are provided in Appendix C.
-For the Reflection & Refinement dimension, ground truth is guaranteed since errors are intentionally produced.
-To further assess the robustness of LLM capabilities, SMART employs a suite of rewriting rules to create augmented versions of the dimension-specific tasks.
-The specifics of these rewriting strategies and the results of associated experiments are detailed in Sec.
-4.3 and 4.4.
-The self-validation methodology remains effective for ascertaining the correctness of these newly formulated, augmented questions, even when their answers differ from those of the original problems.
-This capability is crucial for maintaining the integrity of dynamically generated test instances.
-A comprehensive description of the procedures for the self-validation of augmented data is available in Appendix C.
-4 Experiments
-4.1 Experimental Setup
-We evaluate 21 recent open-source and closed-source LLMs using our SMART evaluation framework.
-The open-source models include Phi4 Abdin et al.
-(2024), Gemma3 Team et al.
-(2025), GLM4 GLM et al.
-(2024), Qwen2.5 Team (2024), Qwen3 Yang et al.
-(2025), Llama3 AI@Meta (2024), Llama4 Meta (2025), and Mistral MistralAITeam (2024).
-The closed-source models assessed are GPT-4o Achiam et al.
-(2023), GPT-4.1 OpenAI (2025a), o4-mini, o3 OpenAI (2025b), DeepSeek-V3 Liu et al.
-(2024), DeepSeek-R1 Guo et al.
-(2025), Claude3.5 anthropic (2024), Claude3.7 anthropic (2025), and Gemini2.5 Google (2024).
-The temperature is set to 0 to ensure deterministic outputs.
-We employ an LLM-as-a-judge approach Zheng et al.
-(2023) to evaluate the understanding dimension.
-Tasks for the other dimensions primarily utilize accuracy-based metrics.
-For all dimension-specific tasks, we adopt a three-shot prompting strategy.
-Detailed descriptions of the evaluation metrics and prompt designs for each of the four SMART dimensions are provided in Appendix D.2.
-Model Final Answer Understanding Reasoning Arithmetic Reflection&Refinement Reflection Refinement
-ACC@Fi LLM@Un ACC@Re ACC@Ar ACC@RR ACC@R-t ACC@R-m
+This evaluation is conducted through a numeric calculation task that presents problems exclusively in terms of numerical values and variables, requiring the execution of fundamental arithmetic operations, as illustrated in Fig.
+2 Dimension 3.
+This experimental setup isolates arithmetic skills from other cognitive demands—such as language comprehension or complex reasoning—thereby enabling a focused and precise evaluation of the model’s arithmetic capabilities.
+Reflection & Refinement The fourth dimension of SMART is Reflection & Refinement, which assesses the LLM’s capacity for self-critique and iterative improvement.
+As illustrated in Fig.
+2, Dimension 4, the model is presented with a question and its chain-of-thought (CoT) solution, and is tasked with identifying potential errors in CoT.
+It then revise the errors and produce an corrected CoT.
+A key design feature is that if the LLM fails to detect all errors in the initial solution, it is not allowed to proceed to the refinement stage for that instance.
+Data Curation
+We adopt a self-generating approach to construct SMART-Bench from a set of seed questions. Tab. 1 provides a summary of SMART-Bench task, ground truth, and evaluation for each dimension in SMART.
+Seed Question Collection
+We begin by collecting a diverse set of 2,000 seed questions from seven widely-used mathematical problem datasets: GSM8K (Cobbe et al.
+2021), SVAMP (Patel, Bhattamishra, and Goyal 2021), ASDiv (Miao, Liang, and Su 2020), AQuA (Ling et al.
+2017), MAWPS (Koncel-Kedziorski et al.
+2016), MATH (Hendrycks et al.
+2021), and the AIME 2024 (Huggingface 2024).
+These seed questions serve as the foundation for generating dimension-specific evaluation tasks.
+Test Case Generation
+Each dimension in SMART-Bench is designed with a distinct test case tailored to evaluate specific aspects of LLMs.
+For the Understanding and Reasoning dimensions, the input questions are identical to the original seed problems, but each targets to generate different outputs as shown in Tab.1.
+For the Understanding task, we use structured information extracted by GPT-4.1 and further verified by human annotators as the ground truth.
+Directly converting a seed question into an arithmetic notation problem is challenging, as it requires simplifying complex natural language into structured mathematical operations while preserving logical relationships between variables.
+Therefore, for the Arithmetic dimension, we adopt a two-stage process: seed questions are first formalized into SMT-LIB representations using GPT-4.1 to capture the core logic, and these formal expressions are then translated into arithmetic problems, also with GPT-4.1.
+For the Reflection dimension, the input consists of the original question paired with a chain-of-thought (CoT) solution containing deliberately injected errors.
+We define six categories of errors: arithmetic inaccuracies, omitted steps, hallucinated content, logical disorder, redundancy, and operator misuse.
+All errors are generated based on predefined rules to simulate flawed reasoning processes.
+Detailed descriptions of these CoT errors are provided in the Appendix.
+The Refinement task takes the same input as Reflection task but ask to generate the corrected CoT.
+Quality Control
+As shown in Tab.
+1, only the input questions for the Arithmetic task and the ground truth for the Understanding task are generated by LLMs and thus require additional verification.
+To ensure the quality of SMART-Bench, we implement a self-validating mechanism comprising both neuro-symbolic and human verification procedures.
+This mechanism automatically identifies and filters out low-quality samples.
+If a sample fails the validation process, it is iteratively regenerated until it satisfies the required quality standards.
+Neuro-symbolic verification
+We employ neuro-symbolic verification to ensure the correctness of the SMT-LIB expressions used in the Arithmetic dimension.
+Directly comparing generated SMT-LIB with ground-truth expressions is challenging.
+Instead, we leverage the Z3 solver (De Moura and Bjørner 2008) to automatically compute the result of a symbolic formula and compare it against the ground-truth answer from the original seed question, as a correct SMT-LIB expression should yield the same answer as its seed question.
+This generation–validation process is repeated until the SMT-LIB expression produces the correct answer.
+Model Understanding Reasoning Arithmetic $R$ Reflection Refinement Final Answer All-Pass Score
+LLM@Un ACC@Re ACC@Ar ACC@$R$ ACC@R-t ACC@R-m ACC@Fi ACC@All
 Open-source models
-Phi4-14B 89.55 93.17 64.12 96.45 22.06 26.75 82.34
-Gemma3-27B 54.62 94.47 40.81 39.94 24.23 32.55 74.35
-GLM4-32B 54.13 94.39 55.92 36.62 22.85 28.50 88.57
-Qwen2.5-72B 41.05 95.21 62.53 35.55 33.85 35.75 94.69
-Qwen3-32B 93.75 93.19 86.14 98.50 54.75 56.70 96.52
-Qwen3-30B-A3B 92.85 93.29 85.20 98.15 43.89 58.25 75.36
-Llama3.1-8B 45.15 88.92 9.05 62.15 3.30 8.80 37.46
-Llama3.3-70B 88.95 95.04 49.26 94.72 40.30 41.75 96.53
-Llama4-Scout-17B-16E 91.24 94.48 62.65 97.11 29.04 49.5 58.67
-Mistral-Small-3.1-24B 77.55 94.48 28.96 49.55 24.55 36.65 66.98
-Mistral-Large-123B 62.25 94.75 48.95 43.74 43.12 50.61 81.62
+Phi4-14B 93.41 64.12 96.45 22.06 26.75 82.34 89.55 20.77
+Gemma3-27B 94.31 40.81 39.94 24.23 32.55 74.35 54.62 11.96
+GLM4-32B 94.11 55.92 36.62 22.85 28.50 88.57 54.13 11.56
+Qwen2.5-72B 95.86 62.53 35.55 33.85 35.75 94.69 41.05 13.95
+Qwen3-32B 93.32 86.14 98.50 54.75 56.70 96.52 93.75 50.72
+Qwen3-30B-A3B 93.14 85.20 98.15 43.89 58.25 75.36 92.85 44.11
+Llama3.1-8B 88.58 9.05 62.15 3.30 8.80 37.46 45.15 2.74
+Llama3.3-70B 94.74 49.26 94.72 40.30 41.75 96.53 88.95 37.25
+Llama4-Scout-17B-16E 94.85 62.65 97.11 29.04 49.5 58.67 91.24 29.54
+Mistral-Small-3.1-24B 94.65 28.96 49.55 24.55 36.65 66.98 77.55 16.87
+Mistral-Large-123B 94.85 48.95 43.74 43.12 50.61 81.62 62.25 21.85
 Closed-source models
-GPT-4o 57.65 95.32 64.45 45.64 42.60 45.30 94.04
-GPT-4.1 59.31 95.41 73.29 45.71 59.90 63.80 93.89
-o4-mini 93.75 95.22 90.72 98.25 55.05 64.61 85.22
-o3 91.44 95.17 89.12 98.45 71.51 78.62 90.97
-DeepSeek-V3 93.43 94.92 65.25 98.35 68.65 71.10 96.55
-DeepSeek-R1 97.92 94.88 93.61 98.15 62.50 69.80 89.54
-Claude3.5-Sonnet 92.61 94.69 80.85 97.61 33.05 65.85 50.19
-Claude3.7-Sonnet 93.53 95.09 74.18 98.11 57.43 66.35 86.51
-Gemini2.5-Flash-Preview 93.33 94.25 89.72 98.12 35.85 50.35 71.20
-Gemini2.5-Pro-Preview 94.15 94.23 91.45 97.95 54.80 76.90 71.26
+GPT-4o 95.22 64.45 45.64 42.60 45.30 94.04 57.65 24.87
+GPT-4.1 95.30 73.29 45.71 59.90 63.80 93.89 59.31 31.74
+o4-mini 95.25 90.72 98.25 55.05 64.61 85.22 93.75 56.23
+o3 95.02 89.12 98.45 71.51 78.62 90.97 91.44 64.87
+DeepSeek-V3 95.12 65.25 98.35 68.65 71.10 96.55 93.43 45.74
+DeepSeek-R1 94.92 93.61 98.15 62.50 69.80 89.54 97.92 57.14
+Claude3.5-Sonnet 94.75 80.85 97.61 33.05 65.85 50.19 92.61 33.04
+Claude3.7-Sonnet 94.85 74.18 98.11 57.43 66.35 86.51 93.53 57.91
+Gemini2.5-Flash-Preview 94.69 89.72 98.12 35.85 50.35 71.20 93.33 38.69
+Gemini2.5-Pro-Preview 94.86 91.45 97.95 54.80 76.90 71.26 94.15 56.25
 Table 2: The performance of open and closed-source models on SMART-Bench.
-4.2 Performance on the SMART-Bench
-The performance of the 21 evaluated LLMs on SMART-Bench is detailed in Tab. 1, which presents scores for final answer accuracy alongside metrics for each of the four dimensions.
-Dimension specific-task results.
-Our results indicate that the evaluated models generally demonstrate a strong capacity for problem understanding, with most LLM@Un scores exceeding 90%.
+Human verification
+The generation of arithmetic notation problems from SMT-LIB representations, as well as the extraction of structured information as ground truth for the Understanding dimension, are both performed using GPT-4.1 and cannot be validated by the neuro-symbolic system.
+Thus, we follow the human verification protocol proposed by Chen et al.
+(2024) to ensure the reliability of these generated data.
+Specifically, a randomly selected 10% subset of the data is manually reviewed by human annotators.
+If this sample does not meet quality standards, the entire set is regenerated and re-verified until it passes inspection.
+As a result, SMART-Bench comprises 10,000 test instances, including 2,000 original seed questions and 8,000 carefully curated, task-specific variants.
+Through a combination of neuro-symbolic and human verification, we ensure that each instance meets quality standards.
+Experiments
+Models
+We evaluate 21 recent open- and closed-source LLMs using our SMART framework, covering both general-purpose and reasoning-specialized models.
+To ensure deterministic outputs, the temperature is set to 0.
+For each dimension-specific task, we employ a three-shot prompting strategy.
+Figure 3: The performance across the varying difficulty settings for each SMART dimension.
+Evaluation Metrics
+We employ a combination of evaluators for SMART, including LLM-as-a-Judge (Zheng et al.
+2023), the Z3 solver, and rule matching, as summarized in Tab.
+1.
+Additionally, we introduce a new metric, All-Pass Score, which reflects a model’s overall performance across the full problem-solving process.
+For the Arithmetic, Reflection and Refinement tasks, we employ a rule-based approach to extract the final answer from the model output and compare it against the ground truth to compute the accuracy-based metric.
+For example, the accuracy for the Arithmetic task (denoted as ACC@Ar) is computed as follows:
+$ACC@Ar$ $=$ $\frac{N_{correct\_answers\_sample}}{N_{total\_sample}}$ (1)
+Metrics such as ACC@An (Final answer accuracy), ACC@Re (Reasoning task), ACC@R-t (Reflection task), ACC@R-m (Refinement task), and ACC@RR (R&R task) are computed using the same formulation.
+For Reasoning task, we do not evaluate the correctness of the generated SMT-LIB expressions, since multiple logically equivalent solutions can exist for a problem.
+Instead, we validate these symbolic expressions by executing them with the Z3 solver and comparing the solver’s output with the ground-truth answer from the original seed question.
+For the Refinement task, we GPT-4.1 to generate a final answer based upon the refined COT.
+The the refined COT is correct if the GPT-4.1-generated response matches the ground truth.
+For the Understanding dimension, we adopt the LLM-as-a-Judge evaluation approach, introducing the metric LLM@Un to assess the quality of generated structured contexts.
+To mitigate potential preference bias (Li et al.
+) inherent in LLM-based judging, we independently employ GPT-4.1 and DeepSeek-V3 as two judging models.
+Each judge evaluates the semantic similarity between the model-generated context and the reference ground truth, assigning a similarity score ranging from 1 to 100.
+The final LLM@Un score is computed by averaging the two judges’ scores:
+$LLM@Un$ $=$ $\frac{LLM@Un_{GPT}+LLM@Un_{DS}}{{2}}$ (2)
+The All-Pass Score is an integrated metric (ACC@ALL) that combines performance across all evaluation dimensions.
+Specifically, a model achieves an All-Pass success if it simultaneously meets the following criteria: (1) obtaining a score of at least 90% on the Understanding task; (2) correctly solving the Reasoning task; (3) correctly solving the Arithmetic task; and (4) successfully completing the entire R&R task.
+This stringent criterion ensures that the All-Pass Score accurately reflects comprehensive, end-to-end mathematical problem-solving capability.
+Model
+Perturbation Final Answer Understanding Reasoning Arithmetic R&R
+ACC@An PD LLM@Un PD ACC@Reason PD ACC@Ar PD ACC@$R$ PD
+GPT-4.1
+Seed question 59.31 / 95.41 / 73.29 / 45.71 / 59.90 /
+[…]
+[…]
+[…]
+Claude-3.7
+Seed question 93.53 / 95.09 / 74.18 / 98.11 / 57.43 /
+[…]
+[…]
+[…]
+Table 3: The performance degradation of evaluation dimensions when three types of perturbations are added to the seed questions.
+PD refers to the performance drop.
+The most affected dimension in each case is highlighted in bold.
+Method
+Refined
+Dimension
+ACC@An
+ACC@An
+with self-refine
+Llama3.1-8B Reasoning 45.15 43.85 ($\downarrow 1.30$)
+Mistral-Small-24B Reasoning 77.55 78.15 ($\uparrow 0.61$)
+GLM-24B Arithmetic 54.13 56.28 ($\uparrow 2.15$)
+Qwen2.5-72B Arithmetic 41.05 52.82 ($\uparrow 11.77$)
+Table 4: Self-Refine prompting on specific dimension.
+Performance on the SMART-Bench
+The performance of the 21 evaluated LLMs on SMART-Bench is detailed in Tab. 2, which reports the scores for all evaluation dimensions and final answer accuracy.
+Dimension specific-task results
+Our results indicate that the LLMs generally demonstrate a strong capacity for problem understanding, with most LLM@Un scores exceeding 90%.
 This suggests a general proficiency in grasping relevant information and interpreting problem statements.
-However, significant performance disparities emerge in the reasoning dimension, where ACC@Re scores range widely from 9.05% to 91.45%.
+However, significant performance disparities emerge in the Reasoning dimension, where ACC@Re scores range widely from 9.05% to 93.61%.
 A similar divergence is observed in the Reflection task with the highest score (78.62%) being nearly nine times greater than the lowest (8.80%).
-These findings suggest that symbolic reasoning and error reflection capabilities represent critical bottlenecks, particularly for lower-performing models.
-In contrast, the arithmetic and refinement tasks appear relatively more tractable for most models, with top-performing LLMs achieving near-perfect scores.
-For instance, o3 attains an ACC@Ar of 98.45%, and DeepSeek-V3 achieves an ACC@R-m of 96.55%, underscoring their strengths in computational and corrective abilities.
-Granular insights.
-Furthermore, the SMART-Bench framework reveals nuanced performance differences among LLMs that final answer accuracy can obscure.
+These findings suggest that symbolic reasoning and error reflection capabilities represent critical bottlenecks, particularly for smaller models.
+In contrast, the Arithmetic and Refinement tasks appear relatively less challenging, with leading LLMs achieving near-perfect performance.
+For example, o3 attains an ACC@Ar of 98.45%, while DeepSeek-V3 reaches an ACC@R-m of 96.55%, demonstrating their strength in computational and corrective abilities.
+Granular insights
+The SMART-Bench framework uncovers nuanced performance differences among LLMs that are obscured by final answer metrics alone.
 For example, while o4-mini and Claude3.7-Sonnet exhibit similar final answer accuracies (93.75% and 93.53%, respectively), o4-mini demonstrates markedly higher proficiency in the Reasoning dimension (90.72%) compared to Claude3.7-Sonnet (74.18%).
-A similar pattern is observed when comparing o4-mini with DeepSeek-V3, demonstrating SMART-Bench’s capacity to reveal fine-grained performance gaps beyond final accuracy metrics.
-Additionally, consider o4-mini and GPT-4.1: they achieve comparable scores in the understanding dimension and the reflection & refinement dimension.
-However, o4-mini’s final answer accuracy (93.75%) substantially surpasses that of GPT-4.1 (59.31%).
+A similar trend is observed when comparing o4-mini and DeepSeek-V3, further illustrating SMART-Bench’s ability to reveal fine-grained gaps that traditional outcome-based metrics miss.
+Additionally, although o4-mini and GPT-4.1 perform similarly on the Understanding and Reflection & Refinement dimensions, o4-mini’s final answer accuracy (93.75%) is markedly higher than GPT-4.1’s (59.31%).
 Our framework attributes this disparity primarily to GPT-4.1’s lower Reasoning (73.29% vs.
@@
 Thus, the SMART framework facilitates a deeper analysis and interpretation of the underlying causes for performance differences.
-Data contamination.
-Theoretically, LLMs should achieve higher accuracy on the arithmetic task compared to the final answer accuracy on the more complex seed questions, as the arithmetic tasks present simplified numerical computations in mathematical notation.
-However, our results reveal instances where models such as Gemma3-27B, GLM4-32B, GPT-4o, and GPT-4.1 exhibit lower ACC@Ar scores than their ACC@Fi scores.
-This counter-intuitive finding may suggest the presence of data contamination in the training sets of these models, where they might have memorized answers to full problems similar to those in standard benchmarks but struggle with the decontextualized arithmetic components presented in our SMART-Bench.
-Figure 3: The performance across the varying difficulty settings for each SMART dimension.
-4.3 How Does Task Difficulty Impact Different Dimensions of SMART?
-To examine factors affecting model performance across each SMART dimension, we construct new sets of dimension-specific questions with varying difficulty levels and evaluate five leading closed-source LLMs on this dynamic test set.
-Details of the difficulty design are provided in Appendix D.3.
-Fig.
-3 presents performance trends across difficulty levels for each SMART dimension.
-As expected, increased task complexity generally leads to substantial performance degradation, validating the effectiveness of our difficulty calibration.
-Notably, GPT-4.1 and Claude3.7-Sonnet exhibit heightened sensitivity in the reasoning dimension, with ACC@Re scores dropping sharply from  90% to below 60% as reasoning steps increase.
-In contrast, the remaining models sustain ACC@Re scores above 80% even beyond six reasoning steps.
-In the arithmetic dimension, o4-mini maintains robust performance (ACC@Ar > 80%) even with nine-digit numbers (refers to the number of digits in scientific notation instead of the count of operands), whereas GPT-4.1’s accuracy drops below 10%.
-For reflection tasks, introducing only two error types into the CoT results in a sharp decline in detection accuracy across all models.
-None can reliably detect all errors when four or more distinct mistake types are present.
-Model Perturbation Final Answer Understanding Reasoning Arithmetic
-ACC@An PD LLM@Un PD ACC@Reason PD ACC@Ar PD
-GPT-4.1 Seed question 59.31 / 95.41 / 73.29 / 45.71 /
-+ Noise insertion 41.23 18.08$\downarrow$ 88.52 6.89$\downarrow$ 46.54 26.75$\downarrow$ 39.40 6.31$\downarrow$
-+ Adding operation 28.45 30.86$\downarrow$ 91.32 4.09$\downarrow$ 57.53 15.76$\downarrow$ 20.64 25.07$\downarrow$
-+ Numerical variation 24.81 34.50$\downarrow$ 92.83 2.58$\downarrow$ 52.81 20.48$\downarrow$ 18.34 27.37$\downarrow$
-Claude-3.7 Seed question 93.53 / 95.09 / 74.18 / 98.11 /
-+ Noise insertion 87.60 5.93$\downarrow$ 89.87 5.22$\downarrow$ 65.81 8.37$\downarrow$ 91.82 6.29$\downarrow$
-+ Adding operation 76.21 17.32$\downarrow$ 92.34 2.75$\downarrow$ 54.38 19.80$\downarrow$ 67.43 24.39$\downarrow$
-+ Numerical variation 62.74 30.79$\downarrow$ 93.17 1.92$\downarrow$ 58.92 15.26$\downarrow$ 58.29 39.82$\downarrow$
-Table 2: The performance degradation of evaluation dimensions when three types of perturbations are added to the seed questions. PD refers to the performance drop.
-4.4 How Do Fine-grained Dimensions Influence the Performance of Final Answer Accuracy?
-Prior work has demonstrated that LLMs exhibit significant performance degradation in final answer accuracy when evaluated on question variations generated through perturbations Li et al.
-(2024b); Zhu et al.
-(2023); Li et al.
-(2024a).
-However, the specific contributions of different cognitive dimensions to this degradation remain underexplored.
-Therefore, we adapt three perturbation strategies from Li et al.
-(2024b) applying them to both the seed questions and their corresponding dimension-specific variants.
-This allows us to investigate which underlying capabilities are most influential in the observed drop in final answer accuracy under perturbation.
-Tab.
-2 shows that all evaluated dimensions exhibit noticeable performance drops (PD) under perturbations.
+All-Pass Score remains a challenge
+To provide a more challenging and discriminative evaluation, we introduce the All-Pass Score, which measures a model’s ability to simultaneously satisfy the success criteria across all dimensions for each seed question.
+The All-Pass Score effectively distinguishes model performance.
+The highest All-Pass Score achieved is only 64.87% for o3, which is substantially lower than its final answer accuracy of 91.44%.
+This metric offers a much clearer assessment of a model’s mathematical ability and highlights considerable room for further improvement.
+How Does Task Difficulty Impact Different Dimensions of SMART?
+To investigate how task difficulty affects model performance across different SMART dimensions, we construct new sets of dimension-specific questions with varied difficulty levels and evaluate five leading closed-source LLMs on this dynamic test set.
+Task difficulty is manipulated in the following ways: for the Understanding dimension, by varying the number of added irrelevant sentences; for the Reasoning dimension, by grouping questions according to the number of required reasoning steps; for the Arithmetic dimension, by changing the number of digits (referring to digit length in scientific notation, not the number of operands); and for the R&R dimension, by altering the number of mistakes introduced into the CoT.
+It is important to note that modifying the number of digits in arithmetic questions changes the ground-truth answer.
+To ensure correctness, we simultaneously update both the numerical values in the arithmetic questions and their corresponding SMT-LIB representations, subsequently employing the Z3 solver to generate new ground-truth answers.
+For the other dimensions, the ground-truth answers remain unchanged.
+Comprehensive details regarding the difficulty settings are provided in the Appendix.
+As shown in Fig.
+3, increasing task complexity generally leads to notable performance degradation across all dimensions.
+Notably, GPT-4.1 and Claude3.7-Sonnet show pronounced sensitivity in the Reasoning dimension, with ACC@Re scores dropping sharply from approximately 90% to below 60% as the number of reasoning steps increases.
+In contrast, the remaining models maintain ACC@Re scores above 80% even with more than six reasoning steps.
+In the Arithmetic dimension, o4-mini demonstrates robust performance (ACC@Ar >> 80%) even with nine-digit numbers, whereas GPT-4.1’s accuracy falls below 10%.
+For Reflection tasks, introducing just two error types into the CoT results in a steep decline in detection accuracy for all models, with none able to reliably detect all errors when four or more distinct mistake types are present.
+How Do Fine-grained Dimensions Influence the Performance of Final Answer Accuracy?
+Prior work has shown that LLMs experience significant drops in final answer accuracy when evaluated on perturbed versions of questions (Li et al.
+2024b; Zhu et al.
+2023; Li et al.
+2024a).
+However, the underlying reasons for this degradation remain insufficiently explored.
+To address this gap, we adapt three perturbation strategies from (Li et al.
+2024b) and apply them to both the seed questions and their corresponding dimension-specific variants, aiming to identify which dimensions are most susceptible to performance loss under these perturbations.
+As shown in Table 3, all evaluated dimensions exhibit substantial performance drops (PD) under perturbations.
 When noise is introduced, the reasoning dimension is most affected for both GPT-4.1 (26.75%) and Claude 3.7-Sonnet (8.37%).
-In contrast, perturbations involving added operations or numerical modifications result in the largest performance drops in the arithmetic dimension for both models.
-These findings suggest that irrelevant information primarily disrupts reasoning capabilities, while alterations to operations or numerical values predominantly impair arithmetic proficiency.
-Such fine-grained dimension performance drop ultimately contribute to reduced final answer accuracy.
-Figure 4: The confusion matrix of final answer and other dimensions. P means Positive, and N means Negative.
-4.5 Is the Final Answer Accuracy Reliable for Measuring Mathematical Capability?
-Given the impressive performance of LLMs and the potential for data contamination, there is a concern that models might solve problems correctly without possessing genuine underlying mathematical capability.
-To investigate this, we conduct experiments to compute confusion matrices comparing final answer correctness with performance on our SMART dimensions, as illustrated in Fig.
-4.
-We posit that true mathematical capability is more accurately reflected by instances where LLMs correctly solve not only the original seed question but also simultaneously succeed in the corresponding reasoning and arithmetic dimension tasks.
-Across all confusion matrices, the False Negative (FN) values are consistently non-zero.
-This indicates that LLMs can sometimes arrive at correct final answers through heuristic shortcuts or other opaque mechanisms, even when their intermediate reasoning or calculation processes are flawed.
-For example, GPT-4.1 exhibits a notable FN rate of 11.90% in final answer and arithmetic confusion matrix.
-Similarly, Claude3.7-Sonnet shows an FN rate of 22.95% in final answer and reasoning confusion matrix.
-These FN cases represent instances where final answer accuracy overestimates the model’s grasp of the intermediate steps.
-Conversely, False Positive (FP) scores denote cases where a model successfully completes an intermediate task but ultimately yields an incorrect final answer.
-Except for GPT-4.1, most evaluated LLMs exhibit relatively low FP rates across various confusion matrices, indicating that accurate intermediate reasoning generally correlates with correct final outputs.
-True Positive (TP) scores capture instances where a model not only produces the correct final answer but also performs intermediate reasoning and arithmetic correctly.
-We regard this TP metric as a more reliable indicator of genuine mathematical problem-solving ability.
-For high-performing models such as o4-mini, DeepSeek-R1, and Gemini2.5-Pro-Preview, TP scores in the reasoning & arithmetic confusion matrix closely match their ACC@Fi values in Tab.
-1.
-In contrast, GPT-4.1 and Claude3.7-Sonnet exhibit significantly lower TP scores relative to their ACC@Fi, suggesting that their final answer accuracy may overestimate their true reasoning capabilities.
-5 Conclusion
-We introduce the SMART evaluation framework and its accompanying benchmark, SMART-Bench, for systematic assessment of LLMs in mathematical problem solving.
-SMART decomposes this complex ability into four interpretable dimensions, Understanding, Reasoning, Arithmetic, and Reflection & Refinement, enabling more granular evaluation than traditional answer-based metrics.
-A key feature is its self-generating and self-validating mechanism, which iteratively verifies the correctness of generated test data to ensure reliability and scalability.
-Experiments on 21 open- and closed-source LLMs reveal that reasoning and reflection remain primary bottlenecks.
-Moreover, our analysis highlights the limitations of final-answer accuracy as a sole indicator and introduces a new metric to better capture true model capability.
-We hope SMART serves as a foundation for more comprehensive and interpretable evaluations.
-6 Limitation
-While our proposed framework provides a comprehensive evaluation platform, it is important to acknowledge its inherent limitations.
-In particular, although Z3 and SMT-LIB effectively handle linear, integer, and some nonlinear constraints, their problem-solving capabilities are restricted.
-They are unsuitable for highly complex nonlinear problems and certain NP-complete combinatorial tasks.
-To overcome these limitations, future work will investigate using Lean Moura and Ullrich (2021) to formalize and prove complex mathematical theorems involving higher-order logic and intricate proof structures beyond SMT solvers’ scope.
-Appendix A Polya’s Problem-solving Theory
-Polya first propose the four step problem solving method in the book ’How to solve it’ Pólya and Conway (1957) to systematically answer questions.
+Conversely, additional operations or numerical modifications lead to the greatest drops in the Arithmetic dimension.
+These results suggest that irrelevant information primarily undermines reasoning abilities, while changes to operations or numeric values predominantly impact arithmetic proficiency.
+Ultimately, vulnerabilities across all dimensions collectively reduce final answer accuracy.
+Improving LLMs via Self-Refine Prompting on Weak Dimensions
+To enhance the mathematical capabilities of LLMs, we apply self-refinement prompting specifically to the weakest step identified in the problem-solving process.
+The specific prompts are provided in the Appendix.
+As shown in Tab.4, targeting the reasoning or arithmetic dimension leads to notable performance gains for Gemma3-27B, Mistral-Small, and Qwen2.5-72B.
+In contrast, Llama3.1-8B experiences a slight performance drop, likely due to its limited capacity for self-reflection.
+These results demonstrate that the SMART framework is an effective diagnostic tool for pinpointing a model’s weakest dimension, and that targeted intervention on this dimension can improve mathematical performance.
+Conclusion
+We present the SMART evaluation framework and its corresponding benchmark, SMART-Bench, for the systematic assessment of LLMs in mathematical problem solving.
+SMART decomposes this complex ability into four fine-grained dimensions—Understanding, Reasoning, Arithmetic, and Reflection & Refinement—and introduces a novel All-Pass Score metric, enabling more comprehensive evaluation.
+We propose a self-generating and self-validating mechanism that iteratively verifies the correctness of generated test data to construct SMART-Bench.
+Experiments on 21 open- and closed-source LLMs reveal that Reasoning and Reflection remain the primary bottlenecks, and that targeted improvements to weak dimensions can enhance the mathematical capabilities of LLMs.
+We hope that SMART provides a foundation for more comprehensive and interpretable evaluation of large language models in the future.
+Polya’s Problem-solving Theory
+Polya first propose the four step problem solving method in the book ’How to solve it’ (Pólya and Conway 1957) to systematically answer questions.
 According to his theory, it breaks down the process of problem-solving into four steps (Fig.
-1): 1) understanding the problem.
+4): 1) understanding the problem.
 2) make a plan.
@@
   • The reflection & refinement dimension examines the model’s ability to identify errors in the solution process and make corrections to improve the final answer.
-Appendix B SMART-Bench Generation
-B.1 Seed Question Dataset of SMART
+Figure 4: Illustration of the SMART framework for evaluating LLMs.
+According to Polya’s theory, humans solve mathematical problems through four steps: understanding the problem, devising a plan, carrying out the plan, and looking back.
+Based on these steps, we define four evaluation dimensions and corresponding tasks to systematically assess LLMs.
+SMART-Bench Generation
+Seed Question Dataset of SMART
 The foundation of the SMART benchmark is a seed dataset comprising 2000 problem instances.
-These were curated from seven widely-used mathematical problem datasets: GSM8K Cobbe et al.
-(2021), SVAMP Patel et al.
-(2021), ASDiv Miao et al.
-(2020), AQuA Ling et al.
-(2017), MAWPS Koncel-Kedziorski et al.
-(2016), MATH Hendrycks et al.
-(2021), and problems from the AIME 2024 competition Huggingface (2024).
+These were curated from seven widely-used mathematical problem datasets: GSM8K (Cobbe et al.
+2021), SVAMP (Patel, Bhattamishra, and Goyal 2021), ASDiv (Miao, Liang, and Su 2020), AQuA (Ling et al.
+2017), MAWPS (Koncel-Kedziorski et al.
+2016), MATH (Hendrycks et al.
+2021), and problems from the AIME 2024 competition (Huggingface 2024).
 These initial 2,000 seed questions, along with their subsequently generated dimension-specific variations (four per seed question), form the complete SMART-Bench, totaling 10,000 test instances.
@@
 Fig.  6 presents a data sample in the SMART benchmark, which contains the seed question, the extracted context, the SMT-LIB expression, the arithmetic notation question, the CoT and the final answer.
-B.2 Dimension-Specific Testing Data and Task Description
+Dimension-Specific Testing Data and Task Description
 For each seed question, SMART generates distinct variants to evaluate the four targeted problem-solving dimensions.
 The generation and ground truth creation for each dimensional task are described below.
-B.2.1 Understanding
-To generate ground truth for the understanding dimension, we utilize GPT-4o to perform context extraction.
+Understanding
+To generate ground truth for the understanding dimension, we utilize GPT-4.1 to perform context extraction.
 The extracted context is structured into the following components: Problem Scenario (describing the overall context of the problem), Goal (specifying what needs to be solved), Known Quantities (listing explicitly provided numerical values or facts), Unknown Quantities (identifying variables or values to be determined), Relationships and Constraints (detailing connections and limitations between pieces of information), and Irrelevant Information (pinpointing details not pertinent to the solution).
-The prompt employed to guide GPT-4o in extracting this contextual information for ground truth generation is presented in Fig.
+The prompt employed to guide GPT-4.1 in extracting this contextual information for ground truth generation is presented in Fig.
 14.
-To enhance the performance of GPT-4o in generating these contexts, we employed a three-shot prompting strategy.
-To ensure the quality of this generated ground truth, we adopt the human verification and iterative regeneration strategy outlined by Chen et al.
-(2024).
-Specifically, after the initial generation of ground truth contexts, a random 10% sample of these is manually assessed by human annotators.
-If the annotators find less than 80% of the sampled contexts satisfactory, the entire set of contexts is regenerated using a refined approach, and the verification process is repeated.
-B.2.2 Reasoning
+Reasoning
 For the Reasoning dimension, the task does not involve generating new questions or altering the ground-truth answer from the seed question.
@@
 A match indicates a correct reasoning process as captured by the SMT-LIB formula.
-B.2.3 Arithmetic
+Arithmetic
 The arithmetic ability of LLMs is measured by their performance on pure numeric calculation problems that share the same reasoning logic and final answer as the seed question.
@@
 Therefore, to address this, we first generate an SMT-LIB representation of the seed question, which simplifies the reasoning logic among variables.
-Subsequently, we utilize GPT-4o to convert this SMT-LIB representation into the arithmetic notation problem.
+Subsequently, we utilize GPT-4.1 to convert this SMT-LIB representation into the arithmetic notation problem and then manually checked by human annotators.
 The prompt for that process is shown in Fig.
@@
 8.
-B.2.4 Reflection & Refinement
+Reflection & Refinement
 For the Reflection & Refinement dimension, we first generate Chain-of-Thought (CoT) solutions containing deliberate errors.
@@
 For the subsequent refinement sub-task, direct verification of the LLM-corrected CoT is complex.
-Instead, we evaluate the refined CoT by using GPT-4o to generate a final answer based upon it.
-The LLM is considered to have successfully passed the refinement task if this GPT-4o-generated answer matches the ground truth of the original seed question.
+Instead, we evaluate the refined CoT by using GPT-4.1 to generate a final answer based upon it.
+The LLM is considered to have successfully passed the refinement task if this GPT-4.1-generated answer matches the ground truth of the original seed question.
 Figure 5: The number of reasoning step statistics of seed question dataset.
 Figure 6: A data sample in the SMART benchmark.
-Appendix C SMART-Bench Verification
-We adopt the neuro-symbolic system to automatically verify the correctness of generated SMT-LIB and its arithmetic notation formed question.
-This self-validating mechanism can also be use to verify the augmented version of questions which are generated based on the number of operator modification rewriting rules.
-The process is as following:
-  1. 1.
-    Initial SMT-LIB Conversion and Validation: We first attempt to convert the original seed question  into an SMT-LIB formula.
-We employ an SMT solver to compute the solution from this SMT-LIB formula.
-If the solver-derived result does not match the known ground-truth answer of the original seed problem, the SMT-LIB generation process is iteratively refined and re-verified until consistency is achieved.
-This ensures a validated symbolic representation of the original problem.
-  2. 2.
-    Augmentation in SMT-LIB Space: Once a validated SMT-LIB formula for the original problem  is obtained, numerical or operational changes dictated by our rewriting rules are applied directly within this symbolic SMT-LIB space.
-  3. 3.
-    Automated Ground-Truth Generation for Augmented Problems: The modified SMT-LIB formula, now representing an augmented version of the problem, is then solved using the Z3-solver.
-The result provided by the solver serves as the automatically generated and verified ground-truth answer for this new, dynamically created problem instance.
 Figure 7: The prompt for LLMs to convert the seed question to a symbolic expression.
@@
 Figure 13: The prompt for LLMs to refine the anwer of the seed question based on the correct CoT.
-Appendix D Experiments
-D.1 Experiment Setting
+Limitation
+While our proposed framework provides a comprehensive evaluation platform, it is important to acknowledge its inherent limitations.
+In particular, although Z3 and SMT-LIB effectively handle linear, integer, and some nonlinear constraints, their problem-solving capabilities are restricted.
+They are unsuitable for highly complex nonlinear problems and certain NP-complete combinatorial tasks.
+To overcome these limitations, future work will investigate using Lean (Moura and Ullrich 2021) to formalize and prove complex mathematical theorems involving higher-order logic and intricate proof structures beyond SMT solvers’ scope.
+Experiments
+Experiment Setting
+We evaluate 21 recent open-source and closed-source LLMs using our SMART evaluation framework.
+The open-source models include Phi4 (Abdin et al.
+2024), Gemma3 (Team et al.
+2025), GLM4 (GLM et al.
+2024), Qwen2.5 (Team 2024), Qwen3 (Yang et al.
+2025), Llama3 (AI@Meta 2024), Llama4 (Meta 2025), and Mistral (MistralAITeam 2024).
+The closed-source models assessed are GPT-4o (Achiam et al.
+2023), GPT-4.1 (OpenAI 2025a), o4-mini, o3 (OpenAI 2025b), DeepSeek-V3 (Liu et al.
+2024), DeepSeek-R1 (Guo et al.
+2025), Claude3.5 (anthropic 2024), Claude3.7 (anthropic 2025), and Gemini2.5 (Google 2024).
 All experiments were conducted on a Linux server equipped with two NVIDIA H800 GPUs (80GB).
@@
 The Python version used was 3.9.20, and the version of the Transformers package was 4.46.0.
-D.2 Prompt and Evaluation metrics in SMART
-D.2.1 Understanding
+Prompts for Evaluation in SMART
+Understanding
 We propose a context extraction task to measure the capability of LLMs in understanding math problems.
@@
 14.
-We employed an LLM-as-a-judge method Zheng et al.
-(2023) to evaluate the quality of generated context.
-In this method, GPT-4o is used impartially to assess semantic similarity between the background context extracted by the evaluated LLM and the reference ground-truth, assigning a score from 1 to 100.
 Figure 14: The prompt for LLMs to extract context from seed question.
-D.2.2 Reasoning
+Reasoning
 For the reasoning dimension, we introduce a symbolic formalization task to evaluate the symbolic reasoning capability of LLMs.
@@
 Subsequently, the Z3 solver is used to compute the results of the generated SMT-LIB expression.
-The evaluation metric for this task is the accuracy of the SMT-LIB formula results.
-Specifically, the accuracy of generating correct SMT-LIB expressions, determined by the number of correct SMT-LIB (verified by Z3-solver) divided by the total number of problems.
-D.2.3 Arithmetic
+Arithmetic
 For the arithmetic dimension, we introduce a numeric calculation task to assess the arithmetic ability of LLMs.
@@
 The arithmetic problem is designed to have the same answer and reasoning logic as its corresponding seed question.
-The evaluation metrics for this task is accuracy, defined as the number of arithmetic problems solved correctly divided by the total number of problems.
-D.2.4 Reflection & Refinement
+Reflection & Refinement
 For the reflection & refinement dimension, we propose an error correction task that requires LLMs to detect mistakes in the chain-of-thought (CoT) of seed questions, correct these mistakes, and generate a new answer for the seed question.
@@
 If LLMs successfully detect all mistakes and generate the correct final answer based on the corrected CoT, we consider the model to have passed the error correction task.
-The evaluation metrics for the reflection task is accuracy.
-The accuracy in identifying all errors in the chain-of-thought (CoT) solution, computed as the number of problems where all CoT errors are correctly identified divided by the total number of problems.
-The accuracy of refining the CoT after error identification, measured as the number of problems correctly answered based on the refined CoT divided by the number of problems with correctly identified CoT errors.
-The combined accuracy for both reflection and refinement tasks, calculated as the number of problems correctly answered after CoT refinement divided by the total number of problems.
-D.3 Difficulty setting for Dimension-specific Task
+The Self-Refine Prompting
+We apply self-refinement prompting specifically to the weakest step identified in the problem-solving process to improve the mathematic capability of LLMs.
+The prompts used for self-refinement on the reasoning and arithmetic dimensions are shown in Fig.15 and Fig.16.
+Figure 15: The prompt for LLMs to self-refine the Reasoning dimension.
+Figure 16: The prompt for LLMs to self-refine the Arithmetic dimension.
+Difficulty setting for Dimension-specific Task
 To create dimension-specific questions with varying difficulty levels, we employ the following strategies:
@@
 The ground truth for these modified questions is updated by incorporating these noise sentences into the ’Irrelevant Information’ category of the context extraction template.
-For the reasoning dimension, question complexity is defined by the number of distinct mathematical operations (e.g., +,,×,÷$+,-,\times,\div,mod$) required to formulate the solution.
+For the reasoning dimension, question complexity is defined by the number of distinct mathematical operations (e.g., $\times,\div,mod$) required to formulate the solution.
 Problems are then categorized into multiple difficulty levels based on this operational count.
@@
 We randomly introduce a varying number of distinct error types (from the categories listed in Fig.
-18) into the CoT to create different levels of challenge for the error detection and correction tasks.
+20) into the CoT to create different levels of challenge for the error detection and correction tasks.
 Each seed question undergoes a two-stage transformation process.
@@
 This hierarchical, two-phase generation framework enhances reliability and scalability by enabling comprehensive and granular evaluation while mitigating risks of overfitting and data contamination.
-D.4 Examples for question with different difficulty setting
+Examples for question with different difficulty setting
 Fig.17 shows examples of different difficulty settings for the understanding dimension evaluation.
@@
 Figure 20: Example of CoT with different errors.
-D.5 Zero-shot vs. three-shot
+Zero-shot vs. three-shot
 We conducted additional comparative experiments evaluating model performance with and without three-shot examples across all stages.
@@
 Table 5: Zero-shot and three-shot prompt engineer strategy comparison in SMART.
+Ablation Study for Reasoning Task
+LLMs demonstrate strong capabilities in translating natural language into formal language, and the formal translation process itself is not the primary cause of poor performance on the Reasoning task.
+To verify this, we conducted an ablation study (’CoT-to-SMT-LIB’) where models converted correct natural language CoT into SMT-LIB formulas in Tab.6.
+Accuracy was verified via Z3 execution against ground-truth answers.
+All evaluated models achieved over 90% accuracy, indicating that the main challenge lies in generating correct reasoning paths—not in the formal translation.
+This supports our claim that the Reasoning dimension effectively captures a model’s core mathematical Reasoning ability.
+Method CoT to SMT-LIB Question to SMT-LIB
+GPT-4.1 92.58 73.29
+O4-mini 95.68 90.72
+DeepSeek-R1 97.54 93.61
+Claude3.7-Sonnet 96.25 74.18
+Gemini2.5-Pro 97.24 91.45
+Table 6: Accuracy for transferring CoT and Seed queston to SMT-LIB formula